### v1

# Minimal operations over permutation groups

Abstract
We classify the possible types of minimal operations above an arbitrary permutation group. Above the trivial group, a theorem of Rosenberg yields that there are five types of minimal operations. We show that above any non-trivial permutation group there are at most four such types. Indeed, except above Boolean groups acting freely on a set, there are only three. In particular, this is the case for oligomorphic permutation groups, for which we improve a result of Bodirsky and Chen by showing one of the types in their classification does not exist. Building on these results, we answer three questions of Bodirsky that were previously open.
Keywords: Minimal operation, clone, local clone, permutation group, oligomorphic permutation group, constraint satisfaction problem, homogeneous structure.
MSC classes: 08A40, 03C05, 08A05, 08A70, 68Q25, 03C15, 20B05, 20B07.
1 Introduction
1.1 Motivation
One of the aims of universal algebra is classifying algebras up to term equivalence [szendrei1986clones]: term equivalence gives a language-independent way of studying most important properties of an algebra such as subalgebras, congruences, endomorphisms, etc. Indeed, much of universal algebra focuses on obtaining structural information about an algebra from the existence of a given kind of term operation. This perspective naturally gives rise to the notion of a function clone: a set of operations from some finite power of the domain into itself closed under composition and projections (Definition 1.3). Algebras up to term equivalence correspond to clones: we can associate to each algebra the clone of its term operations, term equivalent algebras give rise to the same clone, and every clone is associated to the algebra with function symbols for all of its operations.
The clones on a given domain $B$ form a lattice with respect to inclusion. Given a transformation monoid $\mathcal{T}$ on $B$, the clones whose unary operations equal $\mathcal{T}$ form an interval in this lattice [szendrei1986clones, Proposition 3.1], called the monoidal interval of $\mathcal{T}$, with the clone $\langle\mathcal{T}\rangle$ generated by the transformation monoid and projections being the smallest element: it consists of all functions obtained from functions in $\mathcal{T}$ by adding dummy variables. Since monoidal intervals partition the lattice of clones, the task of classifying the clones on a given domain can be split into their study. Universal algebra distinguishes itself from group and semigroup theory by its focus on higher arity operations. This might be one reason why most research has been on idempotent algebras and clones, i.e. the interval lying above the trivial monoid, where all unary operations in the clones are trivial. Moving to monoidal intervals for larger monoids $\mathcal{T}$, one would expect the interval to shrink since the larger the monoid, the easier it is for a given operation to generate another modulo $\mathcal{T}$, and semigroup theory begins playing a stronger role in the picture [krokhin1995monoid, PINSKER200859]. Indeed, at the other end of the spectrum, the monoidal intevals above the full monoid $\mathcal{T}_{B}$ and above the group of all permutations $\mathcal{S}_{B}$ are well-understood [burle1967classes, Haddad_Rosenberg_1994, machida2007minimal, PINSKER200859].
It is natural to ask what is the minimal amount of structure that can be found in a given monoidal interval which consists of more than one element. This gives rise to the definition of a minimal clone above $\langle\mathcal{T}\rangle$: a clone $\mathcal{C}\supsetneq\langle\mathcal{T}\rangle$ with no other clone sandwiched in between them. It is easy to see that any minimal clone above $\langle\mathcal{T}\rangle$ is generated by a single operation together with $\langle\mathcal{T}\rangle$, which, when chosen to be of minimal arity, is called a minimal operation above $\langle\mathcal{T}\rangle$. Any unary operation not in $\mathcal{T}$ generates together with $\langle\mathcal{T}\rangle$ a clone outside the monoidal interval of $\mathcal{T}$ all whose operations are essentially unary, in the sense that they depend on a single variable (Definition 1.13). If a minimal operation above $\langle\mathcal{T}\rangle$ is of arity $>1$, it has to be essential, i.e. not essentially unary, and the clone it generates together with $\langle\mathcal{T}\rangle$ lies in the monoidal interval of $\mathcal{T}$. Hence, minimal elements in the monoidal interval above $\langle\mathcal{T}\rangle$ correspond to minimal clones above $\langle\mathcal{T}\rangle$ whose minimal operation is non-unary. Over a finite set, every clone $\mathcal{D}\supsetneq{\langle\mathcal{T}\rangle}$ contains a minimal clone above ${\langle\mathcal{T}\rangle}$. This is frequently false over an infinite set (cf. [csakany2005minimal]).
The question of understanding the minimal clones above the clone $\langle\mathrm{Id}\rangle$ generated by the trivial monoid, i.e. the clone $\mathcal{P}_{B}$ consisting of all projections on $B$, has a long history in universal algebra. A classical theorem of Rosenberg [fivetypes] shows that over a finite set there are only five types of minimal operations:
Theorem 1.1 (Five Types Theorem [fivetypes]).
Let $B$ be finite and let $f$ be a minimal operation above $\langle\mathrm{Id}\rangle$. Then $f$ is of one of the following types:
  1. 1.
    a unary operation;
  2. 2.
    a binary operation;
  3. 3.
    a ternary majority operation;
  4. 4.
    a ternary minority operation of the form $x+y+z$ in a Boolean group $(B,+)$;
  5. 5.
    a $k$-ary semiprojection for some $k\geq 3$.
We describe these operations in Definition 1.12. Whilst splitting the possible minimal operations into five types, Rosenberg does not give necessary and sufficient conditions for an operation of these types to be minimal except for the cases of a unary operation and of a minority. Hence, current research, e.g. [brady2023coarse, machida2024orderly], focuses on describing in more detail the possible behaviours of the remaining three cases. From Post [postlattice] and Csákány [csakany1983all], whose work precedes that of Rosenberg, we know all of the minimal clones on a two-element and a three-element domain. However, even the problem of describing all minimal clones on a four-element domain is still open. We refer the reader to [szendrei2024ivo] for a recent review of this topic, which discusses progress of [brady2023coarse] on the case of binary operations. Other review papers on minimal operations are [quackenbush1995survey] and [csakany2005minimal]. In a sense, a full classification of the minimal clones on a finite domain seems to be harder than that of the maximal clones, for which we have a complete description by Rosenberg [rosenberg1965structure, rosenberg1970uber].
It has been observed in [BodCSP] that basically following the same proof, a weakening of Theorem 1.1 which still features five types of minimal operations can be obtained for arbitrary transformation monoids which do not contain any constant operations (Theorem 1.21). Over non-trivial permutation groups, one might expect stronger statements to hold. This is, in particular, suggested both by a long history of results showing that the monoidal interval of many permutation groups collapses into a single element [Haddad_Rosenberg_1994, palfy1984unary, palfy1982contributions, kearnes2001collapsing, krokhin1995monoid], and by a result of Bodirsky and Chen (Theorem 1.2 below) who obtained a stronger classification with only four types above certain infinite permutation groups. Function clones whose unary part forms a permutation group are of particular importance since on the one hand, they correspond to relational structures which are cores (in the sense of [hell1992core, bodirsky2007cores], see below), and on the other, they promise connections between group theory and universal algebra.
We investigate the problem of classifying minimal operations above permutation groups both over a finite and an infinite domain. Our main interest lies in clones that arise as the higher arity symmetries (i.e. polymorphisms) of some relational structure. Analogously to the case of groups, where, over arbitrary domains, automorphism groups of relational structures correspond to those permutation groups which are closed in the pointwise convergence topology, polymorphism clones of relational structures correspond to topologically closed clones. Hence, we will focus on clones closed with respect to the pointwise convergence topology (Definition 1.4), also known as locally closed clones. This topology trivialises on a finite domain. Consequently, for $G$ a permutation group acting on $B$, we investigate minimality above $\overline{\langle G\rangle}$ in the lattice of closed clones (Definitions 1.8 and 1.9), where $\overline{\langle G\rangle}$ is the closed clone generated by $G$: it consists of essentially unary functions arising by the addition of dummy variables to a unary function which on each finite subset of $B$ agrees with some element of $G$. We remark that focusing on closed clones only increases the generality of our results since the notion of generation it gives rise to is more complex than generation by composition and projections. In particular, our methods also classify minimal operations above permutation groups in the lattice of all clones on a given domain (cf. Remark 1.29). Meanwhile, our focus on closed clones allows us to obtain several results for infinite-domain constraint satisfaction problems (CSPs).
[…]
When studying $\mathrm{CSP}(B)$ in the finite or oligomorphic setting, if the polymorphism clone of $B$, $\mathrm{Pol}(B)$, is essentially unary (and $|B|>1$), then $\mathrm{CSP}(B)$ is NP-hard. Hence, we may focus on the case of $\mathrm{Pol}(B)$ containing some essential operation of arity $>1$. For the purposes of studying $\mathrm{CSP}$s, it is sufficient to consider the case of $B$ being a (model-complete) core, i.e. such that the closure of its automorphism group in the pointwise convergence topology $\overline{\mathrm{Aut}(B)}$ corresponds to its endomorphism monoid $\mathrm{End}(B)$ [bodirsky2007cores]. In general, since $B$ is in a finite language, $\mathrm{Pol}(B)$ contains some operation of arity $>1$ minimal above its unary part [BodChen], which, if $B$ is a core, is $\overline{\langle\mathrm{Aut}(B)\rangle}$.
Indeed, classifying the minimal operations above $\overline{\langle\mathrm{Aut}(B)\rangle}$ is a part of the traditional "bottom-up" approach to a complexity dichotomy for CSPs: first classify the minimal operations above $\overline{\langle\mathrm{Aut}(B)\rangle}$, and then structure the remaining part of the argument by considering which functions may live above different minimal operations and what identities they may satisfy. This was the approach of Schaefer’s original complexity dichotomy for constraint satisfaction problems on a two-element domain [schaefer1978complexity]. Moreover, this strategy was heavily used in an oligomorphic setting [bodirsky2010complexity, bodirsky2015schaefer, kompatscher2018complexity, bodirsky2019constraint], with the minimal operations over the automorphism group of the random graph being classified in [minrandom] for this purpose. Other modern techniques [Smoothapp, mottet2023order, feller2024algebraic, bitter2024generalized], while avoiding this precise approach, also require arguments to find, given the existence of some polymorphism which is not essentially unary, the existence of a binary one.
Motivated by these applications, Bodirsky and Chen [BodChen] previously obtained a Rosenberg-style type classification of minimal operations above oligomorphic permutation groups:
Theorem 1.2 (Four types, oligomorphic case, [BodChen, BodCSP]).
Let $G\curvearrowright B$ be an oligomorphic permutation group on a countably infinite $B$. Let $f$ be minimal above $\overline{\langle G\rangle}$. Then, $f$ is of one of the following four types:
  1. 1.
    a unary operation;
  2. 2.
    a binary operation;
  3. 3.
    a ternary quasi-majority;
  4. 4.
    a $k$-ary quasi-semiprojection for some $3\leq k\leq 2r-s$, where $r$ is the number of $G$-orbitals and $s$ is the number of $G$-orbits.
Again, these operations are defined in Definition 1.12. In this paper, we study the fully general context of $G\curvearrowright B$ being a non-trivial group acting faithfully on $B$, where $B$ is a set (not necessarily finite or countable). Nevertheless, our results will also offer a surprising improvement of Theorem 1.2.
1.2 Main results
1.2.1 The minimal functions theorem
We obtain the following:
Theorem A.
Let $G\curvearrowright B$ be a non-trivial group acting faithfully on $B$ with $s$ many orbits (where $s$ is possibly infinite). Let $f$ be a minimal operation above $\overline{\langle G\rangle}$. Then, $f$ is of one of the following types:
  1. 1.
    a unary operation;
  2. 2.
    a binary operation;
  3. 3.
    a ternary minority operation of the form $\alpha\mathfrak{m}$ for $\alpha\in G$, where
    • $G$ is a Boolean group acting freely on $B$;
    • the operation $\mathfrak{m}$ is a $G$-invariant Boolean Steiner $3$-quasigroup.
  4. 4.
    $f$ is a $k$-ary orbit/semiprojection for $3\leq k\leq s$.
Moreover, in case (3), such operations exist if and only if $s=2^{n}$ for some $n\in\mathbb{N}$ or is infinite. In case (4), if $B$ is finite or $G=\mathrm{Aut}(B)$ for $B$ $\omega$-categorical in a finite relational language, we have that a minimal $k$-ary orbit semiprojection exists for each $2\leq k\leq s$.
The definitions of orbit/semiprojection and $G$-invariant Boolean Steiner $3$-quasigroup are given in Definitions 1.23 and 1.28, and they correspond to strengthenings of being an orbit/semiprojection, and being a minority of the form $x+y+z$ in a Boolean group respectively. We actually obtain stronger results classifying operations satisfying weaker conditions than minimality. We outline these in Subsection 1.4.
A surprising feature of Theorem A is that (quasi-)majorities, item 3 in Theorems 1.1 and 1.2, cannot exist as minimal above any non-trivial permutation group. Moreover, whenever $G$ is not a Boolean group acting freely on $B$ with $2^{n}$ or infinitely many orbits, there are only three possible minimal operations. Since the action of an oligomorphic permutation group is never free, our results improve on Theorem 1.2 from [BodChen] in two ways: (i) we show that minimal quasi-majorities cannot exist, and (ii) for the only remaining type of operations of arity $>2$ (i.e. the quasi-semiprojections), we specify the behaviour on the orbits and bound their arity by the number of orbits rather than the number of orbitals. More generally, no analogue of Theorem A was known even on a finite domain, where it constitutes progress for the study of the monoidal intervals of permutation groups and their collapse (cf. Claim 3.6 in [kearnes2001collapsing]). Moreover, it should have consequences for the study of CSPs both of non-rigid finite structures and of $\omega$-categorical structures.
1.2.2 Binary essential operations and three answers to questions of Bodirsky
Theorem A has various consequences for finding low arity essential operations in polymorphism clones of non-rigid structures, which we explore in Section 3. In particular, some of our original motivation for studying minimal operations above permutation groups comes from the following question of Bodirsky:
Question 1.
(Question 14.2.6 (24) in [BodCSP]) Does every countable $\omega$-categorical model-complete core with an essential polymorphism also have a binary essential polymorphism?
Whilst we give a negative answer to Question 1 in the fully general setting (Corollary 3.8), Theorem A (or more precisely Theorem C) yields a positive answer whenever $G\curvearrowright B$ is transitive (Corollary 3.7). Moreover, since orbit/semiprojections only satisfy trivial (height 1) identities, we get the following surprising result:
Theorem B.
(Theorem 3.14) Let $G\curvearrowright B$ be such that $G$ is not a Boolean group acting freely on $B$. Let $\mathcal{C}$ be a closed clone belonging to the monoidal interval of $\overline{\langle G\rangle}$. Suppose that $\mathcal{C}$ contains no binary essential operation. Then, $\mathcal{C}$ has a uniformly continuous h1-homomorphism to the clone $\mathcal{P}_{\{0,1\}}$ of projections on a two-element set.
In a finite or $\omega$-categorical context, a uniformly continuous h1-homomorphism to $\mathcal{P}_{\{0,1\}}$ is known to imply NP-hardness [Wonderland], meaning that for the purposes of CSPs, in these cases one may assume there is some binary essential operation in $\mathrm{Pol}(B)$ as long as $\mathrm{Aut}(B)$ is non-trivial and not a Boolean group acting freely on $B$.
Previous literature on infinite-domain CSPs also developed techniques to find binary essential polymorphisms given the existence of an essential polymorphism [BodKara, minrandom, Smoothapp]. These rely on $G\curvearrowright B$ having a particular property, the orbital extension property (Definition 3.17), for which there were no known primitive oligomorphic counterexamples. In Subsection 3.4, we give examples of primitive finitely bounded homogeneous structures without the orbital extension property answering negatively another question of Bodirsky:
Question 2.
(Question 14.2.1 (2) in [BodCSP]) Does every primitive oligomorphic permutation group have the orbital extension property?
Hence, our results on finding binary essential operations expand the reach of pre-existing techniques. Our examples come from Cherlin’s recent classification of homogeneous $2$-multitournaments [Cherlinclass].
Often, arguments for the tractability of $\mathrm{CSP}$s rely on finding a binary injective polymorphism [BodKara, bodirsky2010complexity, bodirsky2014minimal, Smoothapp, bodirsky2020complexity]. Since in Theorem B we found general reasons for why binary essential polymorphisms appear in tractable $\mathrm{CSP}$s of $\omega$-categorical structures, it is reasonable to investigate when we can also get a binary injective polymorphism. We conclude the paper with Section 8, where we answer negatively another question of Bodirsky, which asks whether binary injective polymorphisms can always be found when $\mathrm{CSP}(B)$ meets sufficiently strong tractability assumptions:
Question 3.
(Question 14.2.6 (27) in [BodCSP]) Does every $\omega$-categorical structure without algebraicity that can be solved by Datalog also have a binary injective polymorphism?
We give a counterexample to Question 3 which is a reduct of a finitely bounded homogeneous unary structure, i.e. it lies within the scope of above-mentioned dichotomy conjecture of Bodirsky and Pinsker.
Acknowledgements: The first author would like to thank Jakub Rydval and Carl-Fredrik Nyberg Brodda for helpful discussions. He would also like to thank the participants to a reading group on Bodirsky’s book [BodCSP]. Funded by the European Union (ERC, POCOCOP, 101071674). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. This research was funded in whole or in part by the Austrian Science Fund (FWF) [I 5948]. For the purpose of Open Access, the authors have applied a CC BY public copyright licence to any Author Accepted Manuscript (AAM) version arising from this submission.
1.3 Preliminaries
In this subsection we give some basic definitions and results relevant to the rest of the paper. We follow the discussion of [szendrei1986clones] and [BodChen]. We refer the reader to these references for the relevant proofs.
Let $B$ denote a set. For $n\in\mathbb{N}$, $\mathcal{O}^{(n)}$ denotes the set $B^{{B}^{n}}$ of functions $B^{n}\to B$. We write
$\mathcal{O}:=\bigcup_{n\in\mathbb{N}}\mathcal{O}^{(n)}.$
We equip $\mathcal{O}^{(n)}$ with the product topology and then $\mathcal{O}$ with the sum topology, where $B$ was endowed with the discrete topology.
Definition 1.3.
Let $B$ be a set. A function clone over $B$ is a set $\mathcal{C}\subseteq\mathcal{O}$ such that
  • $\mathcal{C}$ contains all projections: for each $1\leq i\leq k\in\mathbb{N}$, $\mathcal{C}$ contains the $k$-ary projection to the $i$th coordinate $\pi_{i}^{k}\in\mathcal{O}^{(k)}$, given by $(x_{1},\dots,x_{k})\mapsto x_{i}$;
  • $\mathcal{C}$ is closed under composition: for all $f\in\mathcal{C}\cap\mathcal{O}^{(n)}$ and all $g_{1},\dots,g_{n}\in\mathcal{C}\cap\mathcal{O}^{(m)}$, $f(g_{1},\dots,g_{n})$, given by
    $(x_{1},\dots,x_{m})\mapsto f(g_{1}(x_{1},\dots,x_{m}),\dots,g_{n}(x_{1},\dots,x_{m})),$
    is in $\mathcal{C}\cap\mathcal{O}^{(m)}$.
Definition 1.4.
Given $\mathcal{S}\subseteq\mathcal{O}$, $\langle\mathcal{S}\rangle$ denotes the smallest function clone containing $\mathcal{S}$. It consists of all functions in $\mathcal{O}$ which can be obtained by compositions and projections from the functions in $\mathcal{S}$. Meanwhile, $\overline{\mathcal{S}}$ denotes the closure of $\mathcal{S}$ in $\mathcal{O}$ with respect to the topology we described. We call a function clone $\mathcal{C}$ closed, when it is closed with respect to such topology. We have that $f\in\overline{\mathcal{S}}$ if for each finite $A\subseteq B$, there is some $g\in\mathcal{S}$ such that $g|_{A}=f|_{A}$. It is easy to see that for $\mathcal{S}\subseteq\mathcal{O}$, $\overline{\langle\mathcal{S}\rangle}$ is the smallest closed function clone containing $\mathcal{S}$. We say that $\mathcal{S}$ locally generates $g$ if $g\in\overline{\langle\mathcal{S}\rangle}$. Frequently, we will just say that $\mathcal{S}$ generates $g$ since we will always work with local generation.
Definition 1.5.
Let $B$ be a relational structure. We say that $f:B^{n}\to B$ is a polymorphism of $B$ if it preserves all relations of $B$: for any such relation $R$,
[…]
We call the set of polymorphisms of $B$ the polymorphism clone of $B$, and denote it by $\mathrm{Pol}(B)$.
Below, we list some basic facts about closed clones. Their proofs can be found in [szendrei1986clones]:
Fact 1.6.
The closed subclones of $\mathcal{O}$ ordered by inclusion form a lattice.
Fact 1.7 ([rosenberg1999locally]).
The closed subclones of $\mathcal{O}$ on a set $B$ correspond to the polymorphism clones of relational structures on $B$.
Definition 1.8.
Let $\mathcal{D}\supsetneq\mathcal{C}$ be closed subclones of $\mathcal{O}$. We say that $\mathcal{D}$ is minimal above $\mathcal{C}$ if there is no closed clone $\mathcal{E}$ such that
$\mathcal{C}\subsetneq\mathcal{E}\subsetneq\mathcal{D}.$
Definition 1.9.
Let $\mathcal{C}$ be a closed subclone of $\mathcal{O}$. We say that an operation $f\in\mathcal{O}\setminus\mathcal{C}$ is minimal above $\mathcal{C}$ if $\overline{\langle C\cup\{f\}\rangle}$ is minimal above $\mathcal{C}$ and every operation in $\overline{\langle C\cup\{f\}\rangle}\setminus\mathcal{C}$ has arity greater than or equal to that of $f$.
Fact 1.10.
Let $\mathcal{C}$ be a closed subclone of $\mathcal{O}$. Then, the closed subclone $\mathcal{D}\subseteq\mathcal{O}$ is minimal above $\mathcal{C}$ if and only if there is some operation $f$ on $B$ which is minimal above $\mathcal{C}$ and such that $\overline{\langle\mathcal{C}\cup\{f\}\rangle}=\mathcal{D}$.
Notation 1.11.
We use the symbol $\approx$’ to denote identities which hold universally. For example, instead of
$\forall x,y\in B\ f(x,y)=f(y,x),$
we write
$f(x,y)\approx f(y,x).$
Definition 1.12.
We define some types of operations relevant to the rest of the paper by the identities that they satisfy:
  • a ternary quasi-majority is a ternary operation $m$ such that
    $m(x,x,y)\approx m(x,y,x)\approx m(y,x,x)\approx m(x,x,x);$
  • a quasi-Malcev operation is a ternary operation $M$ such that
    $M(x,y,y)\approx M(y,y,x)\approx M(x,x,x);$
  • a ternary quasi-minority is a ternary operation $\mathfrak{m}$ such that
    $\mathfrak{m}(x,x,y)\approx\mathfrak{m}(x,y,x)\approx\mathfrak{m}(y,x,x)\approx\mathfrak{m}(y,y,y);$
  • a quasi-semiprojection is a $k$-ary operation $f$ such that there is an $i\in\{1,\dots,k\}$ and a unary operation $g$ such that whenever $(a_{1},\dots,a_{k})$ is a non-injective tuple from $B$,
    $f(a_{1},\dots,a_{k})=g(a_{i}).$
For each of the operations above, we remove the prefix "quasi" when the operation is idempotent, i.e. satisfies $f(x,\dots,x)\approx x$. For example, a ternary majority is a ternary quasi-majority $m$ which also satisfies $m(x,x,x)\approx x$.
Definition 1.13.
We say that a $k$-ary operation $f$ is essentially unary if there is a unary operation $g$ and $1\leq i\leq k$ such that
$f(x_{1},\dots,x_{k})\approx g(x_{i}).$
We say that $f$ is essential if it is not essentially unary. A clone $\mathcal{C}$ is essentially unary if all of its operations are essentially unary.
1.4 Classifying minimal operations via almost minimal operations
Our path towards Theorem A consists in first classifying operations satisfying weaker versions of minimality and then refining the types of operations obtained from this classification as we approach minimality. The main notion that we investigate is that of almost minimality (Definition 1.17), which captures that $f$ does not locally generate anything new of strictly lower arity together with $\overline{\langle G\rangle}$. Already looking at almost minimality, for $G\curvearrowright B$ not being a Boolean group acting freely on $B$, we get just three types of almost minimal operations. In the case of Boolean groups, after classifying almost minimal operations, we need the further refinement of studying strictly almost minimal operations (Definition 5.8): functions such that every operation of lower or equal arity they locally generate with $\overline{\langle G\rangle}$ locally generates them back with $\overline{\langle G\rangle}$ (but an operation of larger arity might not). Finally, once strictly almost minimal operations are understood, we move to minimal operations by adding the requirement that also operations of strictly larger arity which are locally generated by $f$ together with $\overline{\langle G\rangle}$ locally generate $f$ back with $\overline{\langle G\rangle}$.
Notation 1.14.
Let $G$ be a group acting on a set $B$. By $G\curvearrowright B$ we mean that $G$ acts faithfully on $B$. Moreover, throughout this paper, we assume that $G$ is non-trivial.
Definition 1.15.
A Boolean group is a group where every element has order $2$; in other words, an elementary Abelian 2-group.
Definition 1.16.
We say that the action $G\curvearrowright B$ is free if no non-trivial element in $G$ fixes a point of $B$.
1.4.1 Almost minimal operations
Definition 1.17.
Let $G\curvearrowright B$. The $k$-ary operation $f\in\mathcal{O}\setminus\overline{\langle G\rangle}$ is almost minimal if $\overline{\langle G\cup\{f\}\rangle}\cap\mathcal{O}^{(r)}=\overline{\langle G\rangle}\cap\mathcal{O}^{(r)}$ for each $r<k$.
Remark 1.18.
There is no harm in similarly defining a $k$-ary operation being almost minimal above an arbitrary closed clone $\mathcal{D}$ (rather than $\overline{\langle G\rangle}$). Moreover, note that if $f$ is minimal above $\mathcal{D}$, then it is also almost minimal above it.
Remark 1.19.
Unary functions in $\mathcal{O}^{(1)}\setminus\overline{\langle G\rangle}$ are vacuously almost minimal.
Whilst minimal operations do not necessarily exist over an infinite domain $B$ (cf. [csakany2005minimal]), it is easy to see that almost minimal operations exist over any domain $B$ and any permutation group $G\curvearrowright B$:
Lemma 1.20.
Let $G\curvearrowright B$. Suppose that $\mathcal{C}$ is a clone strictly containing $\overline{\langle G\rangle}$. Then, $\mathcal{C}$ contains some almost minimal operation.
Proof.
We prove that if $\mathcal{C}\supseteq\overline{\langle G\rangle}$ has no almost minimal operation, then $\mathcal{C}=\overline{\langle G\rangle}$. Suppose that $f$ is of minimal arity such that $f\in\mathcal{C}\setminus\overline{\langle{G}\rangle}$. By Remark 1.19, $f$ has arity $>1$. Since $f$ is not almost minimal, it generates some $g$ of strictly smaller arity in $\mathcal{C}\setminus\overline{\langle{G}\rangle}$. But this contradicts the choice of $f$ as being of minimal arity. Hence, $\mathcal{C}=\overline{\langle{G}\rangle}$ as desired. ∎
As mentioned above, in this paper we also classify the almost minimal operations above permutation groups. In order to do this, we begin from the observation that the following version of Rosenberg’s five types theorem [fivetypes] in a non-idempotent context given in [BodCSP, Theorem 6.1.42] only uses the assumption of almost minimality (see Remark 1.22 below):
Theorem 1.21 (Five types theorem over essentially unary clones, [BodCSP]).
Let $\mathcal{C}$ be an essentially unary clone without constant operations and let $f$ be minimal above $\mathcal{C}$. Then, up to permuting its variables, $f$ is of one of the following five types:
  1. 1.
    a unary operation;
  2. 2.
    a binary operation;
  3. 3.
    a ternary quasi-majority operation;
  4. 4.
    a quasi-Malcev operation;
  5. 5.
    a $k$-ary quasi-semiprojection for some $k\geq 3$.
Remark 1.22.
Inspecting the proof of Theorem 1.21 in [BodCSP, §6.1.8], one can see that the only property of minimal operations being used is almost minimality. In particular, Theorem 1.21 actually gives a classification into five types of the almost minimal operations above an essentially unary clone (not containing any constant operations). The proof of Theorem 1.21 follows the same structure as the first part of the standard proof of Rosenberg’s Theorem 1.1 (cf. [fivetypes, csakany2005minimal]).
We classify almost minimal operations above permutation groups $G\curvearrowright B$ using Theorem 1.21 for $\mathcal{C}=\overline{\langle G\rangle}$ as a starting point. Firstly, in Lemma 2.1, we prove that for any permutation group $G\curvearrowright B$, a $k$-ary quasi-semiprojection almost minimal above $\overline{\langle G\rangle}$ satisfies the stronger property of being an orbit/semiprojection:
Definition 1.23.
Let $G\curvearrowright B$. We say that the $k$-ary operation $f$ on $B$ is an orbit/semiprojection, if there is an $i\in\{1,\dots,k\}$ and a unary operation $g\in\overline{G}$ such that for any tuple $(a_{1},\dots,a_{k})$ where at least two of the $a_{j}$ lie in the same $G$-orbit,
$f(a_{1},\dots,a_{k})=g(a_{i}).$
Our classification of almost minimal operations above permutation groups splits into three cases. The first, considered in Section 2, is that of $G\curvearrowright B$ not being a Boolean group acting freely on $B$. In this case, we are able to show that neither quasi-Malcev operations nor quasi-majorities from Theorem 1.21 can be almost minimal above $\overline{\langle G\rangle}$, yielding the following:
Theorem C (Theorem 2.7, Three Types Theorem).
Let $G\curvearrowright B$ be such that $G$ is not a Boolean group acting freely on $B$. Let $s$ be the (possibly infinite) number of orbits of $G$ on $B$. Let $f$ be an almost minimal operation above $\overline{\langle G\rangle}$. Then, $f$ is of one of the following types:
  1. 1.
    a unary operation;
  2. 2.
    a binary operation;
  3. 3.
    a $k$-ary orbit/semiprojection for $3\leq k\leq s$.
In Proposition 2.11 we note that for any $G\curvearrowright B$ (including Boolean groups acting freely), almost minimal $k$-ary orbit/semi-projections exist for all $2\leq k\leq s$.
We can also classify the almost minimal functions above the clone locally generated by a Boolean group acting freely on a set. In this context, freeness of $G\curvearrowright B$ yields that $\overline{\langle G\rangle}=\langle G\rangle$ (Remark 4.1). For $|G|>2$, a new type of almost minimal function is possible, which we call a $G$-quasi-minority:
Definition 1.24.
Let $G\curvearrowright B$. We say that a ternary operation $\mathfrak{m}$ is a $\bm{G}$-quasi-minority if for all $\beta\in G$,
$\mathfrak{m}(y,x,\beta x)\approx\mathfrak{m}(x,\beta x,y)\approx\mathfrak{m}(x,y,\beta x)\approx\mathfrak{m}(\beta y,\beta y,\beta y).$ (1)
We say that a $G$-quasi-minority is a $\bm{G}$-minority if it is idempotent.
Remark 1.25.
For a $G$-quasi-minority $\mathfrak{m}$ to make sense as a possible almost minimal function above $\overline{\langle G\rangle}$, we really need $G$ to be a Boolean group acting freely on $B$. Firstly, condition 1 implies that for all $\beta\in G$,
$\mathfrak{m}(x,x,x)\approx\mathfrak{m}(\beta^{-1}x,x,\beta x)\approx\mathfrak{m}(\beta^{2}x,\beta^{2}x,\beta^{2}x).$
Since all elements in $\overline{\langle G\rangle}\cap\mathcal{O}^{(1)}$ are injective functions, if $m$ is almost minimal above $\overline{\langle G\rangle}$, we must have that the map $x\mapsto\mathfrak{m}(x,x,x)$ (which is generated by $\mathfrak{m}$) is injective. This yields that $\beta^{2}=1$ for all $\beta\in G$, and hence $G$ is Boolean. Secondly, if $\alpha\in G$ fixes a point $a\in B$, for all $b\in B$, by condition 1,
$\mathfrak{m}(b,b,b)=\mathfrak{m}(b,a,a)=\mathfrak{m}(b,a,\alpha a)=\mathfrak{m}(\alpha b,\alpha b,\alpha b).$
By injectivity of $x\mapsto\mathfrak{m}(x,x,x)$, we must have that $\alpha b=b$ for all $b\in B$, which, since $G\curvearrowright B$ is faithful, implies that $\alpha$ is the identity in $G$. Hence, $G\curvearrowright B$ must be free.
Conversely, almost minimal $G$-quasi-minorities always exist whenever $G\curvearrowright B$ is a Boolean group acting freely on $B$, as we show in Proposition 4.8.
Thus, for the case of $G\curvearrowright B$ being the free action of a Boolean group of size $>2$, we get the following classification of almost minimal operations:
Theorem D (Theorem 4.7, Boolean case).
Let $G\curvearrowright B$ be a Boolean group acting freely on $B$ with $s$-many orbits (where $s$ is possibly infinite) and $|G|>2$. Let $f$ be an almost minimal operation above $\langle G\rangle$. Then, $f$ is of one of the following types:
  1. 1.
    a unary operation;
  2. 2.
    a binary operation;
  3. 3.
    a ternary $G$-quasi-minority;
  4. 4.
    a $k$-ary orbit/semiprojection for $3\leq k\leq s$.
In the only remaining case of $\mathbb{Z}_{2}$ acting freely on $B$, there are two more types of almost minimal operations; these actually generate each other.
Definition 1.26.
Let $G\curvearrowright B$. Then, we say that $m(x,y,z)$ is an odd majority if $m$ is a quasi-majority such that for all $\gamma\in G\setminus\{1\}$,
$m(y,x,\gamma x)\approx m(x,\gamma x,y)\approx m(x,y,\gamma x)\approx m(y,y,y).$ (2)
Definition 1.27.
Let $G\curvearrowright B$. The ternary operation $M$ is an odd Malcev if $M$ is a quasi-Malcev such that for all $\gamma\in G\setminus\{1\}$,
$M(x,y,x)\approx M(x,x,x)$ (3)
$M(y,\gamma x,x)\approx M(x,\gamma x,y)\approx M(x,x,x)$ (4)
$M(x,y,\gamma x)\approx M(\gamma y,\gamma y,\gamma y).$ (5)
Theorem E (Theorem 6.7).
Let $\mathbb{Z}_{2}\curvearrowright B$ freely with $s$-many orbits (where $s$ is possibly infinite). Let $f$ be an almost minimal operation above ${\langle\mathbb{Z}_{2}\rangle}$. Then, $f$ is of one of the following types:
  1. 1.
    a unary operation;
  2. 2.
    a ternary $G$-quasi-minority;
  3. 3.
    an odd majority;
  4. 4.
    an odd Malcev (up to permutation of variables);
  5. 5.
    a $k$-ary orbit/semiprojection for $2\leq k\leq s$.
Odd majorities and odd Malcev operations above $\mathbb{Z}_{2}$ are in a sense dual of each other, since an odd majority $m(x,y,z)$ always generates an odd Malcev $m(x,\gamma y,z)$ for $\gamma\in\mathbb{Z}_{2}\setminus\{0\}$ and vice versa. In Proposition 6.9 we prove that almost minimal odd majorities and odd Malcev operations always exist over $\langle\mathbb{Z}_{2}\rangle$.
1.4.2 Back to minimality
We can actually prove that there are no minimal odd majorities or odd Malcevs (Corollary 6.11). This allows us to treat uniformly all Boolean groups acting freely on a set when classifying minimal operations above $\langle G\rangle$ (as opposed to almost minimal operations). Moreover, we are able to prove that minimal $G$-quasi-minorities satisfy some particularly nice properties:
Definition 1.28.
A $\bm{G}$-invariant Boolean Steiner $\bm{3}$-quasigroup is an idempotent symmetric ternary minority operation $m$ also satisfying the following conditions:
$m(x,y,m(x,y,z))$ $\approx z\;;$ (SQS)
$m(x,y,m(z,y,w))$ $\approx m(x,z,w)\;;$ (Bool)
for $\text{for all }\alpha,\beta,\gamma\in G,m(\alpha x,\beta y,\gamma z)$ $\approx\alpha\beta\gamma m(x,y,z).$ (Inv)
As we discuss more in Section 5, idempotent symmetric minorities satisfying (SQS) are called Steiner $3$-quasigroups or Steiner Skeins in the literature and have been studied in universal algebra and design theory [quackenbush1975algebraic, armanious1980algebraische, lindner1978steiner]. When they further satisfy (Bool), they are called Boolean Steiner $3$-quasigroups, since they are known to be of the form $m(x,y,z)=x+y+z$ in a Boolean group. The condition (Inv) is novel and adds further constraints on the behaviour of $m$: it must induce another Boolean Steiner $3$-quasigroup on the orbits which is related to $m$ in a particular way. Indeed, we can characterise $G$-invariant Boolean Steiner $3$-quasigroups in much detail, and we do so in Section 5.2. All of this yields the following theorem:
Theorem F (Theorem 6.12, minimal operations, Boolean case).
Let $G\curvearrowright B$ be a Boolean group acting freely on $B$ with $s$-many orbits (where $s$ is possibly infinite). Let $f$ be a minimal operation above $\langle G\rangle$. Then, $f$ is of one of the following types:
  1. 1.
    a unary operation;
  2. 2.
    a binary operation;
  3. 3.
    a ternary minority of the form $\alpha\mathfrak{m}$ where $\mathfrak{m}$ is a $G$-invariant Boolean Steiner $3$-quasigroup and $\alpha\in G$;
  4. 4.
    a $k$-ary orbit/semiprojection for $3\leq k\leq s$.
Furthermore, all $G$-invariant Boolean Steiner $3$-quasigroups on $B$ are minimal, and they exist if and only if $s=2^{n}$ for some $n\in\mathbb{N}$ or is infinite.
Theorems C and F give us a classification of the types of minimal operations above a non-trivial permutation group, with the latter theorem giving necessary and sufficient conditions for the existence of the minimal operation of the minority type. We would like to also have necessary and sufficient conditions for the existence of minimal orbit/semiprojections in a given arity $k$. Above the trivial group, Pálfy proved that on a domain of size $s$ there are minimal semiprojections for each size $2\leq k\leq s$ [palfy1982contributions]. A similar construction also yields necessary and sufficient conditions for a $k$-ary orbit/semiprojection to exist in the settings most relevant to us:
Theorem G.
(Theorem 7.10, Pálfy’s Theorem for orbit/semiprojections) Let $G\curvearrowright B$ with $s$-many orbits with $B$ finite or of the form $\mathrm{Aut}(B)\curvearrowright B$ for $B$ $\omega$-categorical in a finite relational language. Then, for all $2\leq k\leq s$, there is a $k$-ary orbit/semiprojection minimal above $\overline{\langle G\rangle}$.
The various results mentioned in this section are then summarised in Theorem A:
Proof of Theorem A.
Since minimal operations are almost minimal, Theorem C already tells us the best possible classification of minimal operations above $\overline{\langle G\rangle}$ when $G$ is not Boolean acting freely on $B$. The latter case is dealt with in Theorem F. Finally, Theorem 7.10 gives necessary and sufficient conditions for the existence of minimal orbit/semiprojections in a finite or $\omega$-categorical in a finite relational language context. ∎
Remark 1.29 (Minimality above permutation groups in the lattice of all clones).
Our results also yield a classification of minimal and almost minimal clones above the clone $\langle G\rangle$ generated by a permutation group, when considering minimality in the lattice of all clones rather than that of closed clones. In this context, for a set of operations $\mathcal{S}$, we work with the notion of generation given by $\langle\mathcal{S}\rangle$, rather than $\overline{\langle\mathcal{S}\rangle}$ in the definitions of minimal clone (1.8), minimal operation (1.9), and almost minimal operation (1.17). With these (different) notions of minimality, Theorems C, D, E, F are all true when replacing $\overline{\langle G\rangle}$ by the clone $\langle\mathcal{T}\rangle$ generated by any transformation monoid $\mathcal{T}$ such that $G\subseteq\mathcal{T}\subseteq\overline{G}$. This can be seen by noting that at every step where we exclude the existence of a certain type of minimal (or almost minimal) operation, we do so by proving that if an operation $f$ of that type exists, then it generates (just by composition with itself and group elements and without local closure) another operation which does not locally generate it back. For Theorem G, we actually get the better result that, for all $G\curvearrowright B$ with $s$-many orbits and all $1<k\leq s$, there are minimal (in the lattice of all clones) $k$-ary orbit/semiprojections above $\langle G\rangle$ (Corollary 7.8). This yields a version of Theorem A for minimality above $\langle G\rangle$ in the lattice of all clones where the possible types of minimal operations are exactly those described in Theorem A, and with necessary and sufficient conditions for the existence of those of arity $>2$. Firstly, for $G\curvearrowright B$ a Boolean group acting freely on $B$, a $G$-invariant Boolean Steiner $3$-quasigroup is always minimal above $\langle G\rangle$ both in the interval of all clones and in the interval of closed clones above $\langle G\rangle$ (Lemma 5.18). Hence, such an operation exists and is minimal above $\langle G\rangle$ if and only if the number $s$ of $G$-orbits is either $2^{n}$ or infinite. Secondly, as mentioned above, there is a minimal (in the lattice of all clones) $k$-ary orbit/semiprojection above $\langle G\rangle$ if and only if $1<k\leq s$.
2 A three types theorem
The main result of this section is Theorem 2.7, which classifies almost minimal operations above $\overline{\langle G\rangle}$ for $G$ not being a Boolean acting freely on $B$. We use as a starting point Theorem 1.21 and explore in more detail the consequences of almost minimality for each of the types of operations of arity $>2$. We begin with Lemma 2.1, which, with no assumption on $G$, shows that almost minimal quasi-semiprojection are orbit/semiprojections. Then, Lemma 2.2 and 2.3 show that we need $G\curvearrowright B$ to be the free action of a Boolean group in order for almost minimal quasi-Malcev operations to possibly exist. We will see this is a sufficient condition for them to exist in Section 4. Thus, we proceed by considering the case of a quasi-majority and prove that these can also only exist as almost minimal if $G$ is a Boolean group acting freely on $B$ in Lemma 2.6. Indeed, later in Lemma 4.2, we show that almost minimal quasi-majorities only exist for the free action of $\mathbb{Z}_{2}$. Thus, Theorem 2.7 follows from putting together these results.
Lemma 2.1.
Let $G\curvearrowright B$. For $k\geq 3$, let $f$ be a $k$-ary quasi-semiprojection which is almost minimal above $\overline{\langle G\rangle}$. Then, $f$ is an orbit/semiprojection.
Proof.
We may say without loss of generality that there is a unary operation $g$ such that whenever |$|\{a_{1},\dots,a_{k}\}|<k$, then
$f(a_{1},\dots,a_{k})=g(a_{1}).$
Let $\alpha\in G$. For all $i<j$, the function $h_{i,j,\alpha}$ defined by
$(x_{1},\ldots,x_{k})\mapsto f(x_{1},\ldots,x_{j-1},\alpha x_{i},x_{j+1},\ldots,x_{k})$
does not depend on $x_{j}$, hence it depends on only one variable by the almost minimality of $f$. Moreover, setting for any $\ell\notin\{1,j\}$ the variable $x_{\ell}$ equal to $x_{1}$ and applying $h_{i,j,\alpha}$ yields $g(x_{1})$, by the property on $f$ assumed above. Hence, the variable $h_{i,j,\alpha}$ depends on is $x_{1}$, and
$h_{i,j,\alpha}(x_{1},\ldots,x_{k})\approx g(x_{1})\;.$
Whenever $(a_{1},\ldots,a_{k})$ is any tuple where $a_{i},a_{j}$ belong to the same orbit, then $a_{j}=\alpha a_{i}$ for some $\alpha\in G$. By the above, $f$ will return $g(a_{1})$ on this tuple; hence, $f$ is an orbit/semiprojection onto $x_{1}$ as witnessed by $g$. ∎
Lemma 2.2.
Let $G\curvearrowright B$ be such that the action of $G$ on $B$ is not free. Then, no almost minimal function over $\overline{\langle G\rangle}$ can be a quasi-Malcev operation.
Proof.
Since the action of $G$ on $B$ is not free, there is some non-trivial $\alpha\in G$ and distinct $a,b,c\in B$ such that $\alpha(a)=a,\alpha(b)=c$. Suppose by contradiction that $M(x,y,z)$ is almost minimal and quasi-Malcev, and consider $h(x,y)=M(x,\alpha x,y)$. If $h(x,y)$ depends on the first argument, we have that
$M(a,a,a)=h(a,a)=h(a,b)=M(a,a,b)=M(b,b,b)\;,$
contradicting injectivity of $M(x,x,x)\in\overline{\langle G\rangle}$. Similarly, if $h(x,y)$ depends on the second argument,
$M(c,c,c)=M(a,a,c)=h(a,c)=h(b,c)=M(b,c,c)=M(b,b,b)\;,$
again contradicting injectivity of $M(x,x,x)\in\overline{\langle G\rangle}$. Thus, $h(x,y)$ depends on both arguments, contradicting the almost minimality of $M$. ∎
Lemma 2.3.
Let $G\curvearrowright B$. Suppose some element of $G$ has order $\geq 3$. Then, no almost minimal function over $\overline{\langle G\rangle}$ can be a quasi-Malcev operation.
Proof.
Let $\alpha\in G$ have order $\geq 3$. Since $\alpha^{2}\neq 1$, there is $b\in B$ such that $b,\alpha(b)=c,$ and $\alpha^{2}(b)=d$ are all distinct. Suppose that $M(x,y,z)$ is almost minimal and quasi-Malcev. Consider $h(x,y)=M(x,y,\alpha x)$. We have that
$h(c,d)=M(c,d,d)=M(c,c,c)=M(b,b,c)=h(b,b).$
By almost minimality, $h\in\overline{\langle G\rangle}$, and so it equals a unary injective function of either its first or its second argument. However, since $b,c,d$ are distinct, this yields a contradiction. ∎
Lemma 2.4.
Let $G\curvearrowright B$. Suppose that $m(x,y,z)$ is a quasi-majority operation which is almost minimal above $\overline{\langle G\rangle}$. Then, for any $\beta\in G\setminus\{1\}$,
$h(x,y,z):=m(x,\beta y,z)$
is a quasi-Malcev operation such that
$h(x,x,x)\approx m(x,x,x).$ (6)
Proof.
Note that $h$ is also almost minimal above $\overline{\langle G\rangle}$ since $m(x,y,z)=h(x,\beta^{-1}y,z)$. Hence, the binary function given by $l(x,y):=h(x,x,y)$ must be in $\overline{\langle G\rangle}$ and so essentially unary. Note that $l(x,y)$ cannot depend on the first variable. Otherwise,
$m(x,x,x)$ $\approx m(x,\beta x,x)\approx h(x,x,x)\approx l(x,x)$
$\approx l(x,\beta x)$ $\approx h(x,x,\beta x)\approx m(x,\beta x,\beta x)\approx m(\beta x,\beta x,\beta x),$
which contradicts injectivity of $m(x,x,x)\in\overline{\langle G\rangle}$. Hence, $l(x,y)$ depends on the second variable yielding
$h(x,x,y)\approx l(x,y)\approx l(y,y)\approx h(y,y,y)\approx m(y,\beta y,y)\approx m(y,y,y).$
From this equation we obtain that $h$ satisfies 6. By symmetry, the same argument applies for $h(y,x,x)$, yielding that
$h(x,x,y)\approx h(y,y,y)\approx h(y,x,x),$
and so that $h$ is quasi-Malcev. ∎
Remark 2.5.
Note that by symmetry, from Lemma 2.4, if $m$ is a quasi-majority almost minimal above $\overline{\langle G\rangle}$ and $\beta\in G\setminus\{1\},l(x,y,z)=m(x,z,\beta y)$ is also quasi-Malcev and such that $l(x,x,x)\approx m(x,x,x)$.
Lemma 2.6.
Let $G\curvearrowright B$. Suppose no almost minimal function over $\overline{\langle G\rangle}$ is quasi-Malcev. Then, no almost minimal function over $\overline{\langle G\rangle}$ can be a ternary quasi-majority operation.
Proof.
Suppose that $m$ is a ternary almost minimal quasi-majority above $\overline{\langle G\rangle}$. Then, for $\beta\in G\setminus\{1\}$, $m(x,\beta y,z)$ is quasi-Malcev and almost minimal above $\overline{\langle G\rangle}$ by Lemma 2.4. Since we are assuming there are no quasi-Malcev operators almost minimal above $\overline{\langle G\rangle}$, this yields a contradiction. ∎
Theorem 2.7 (Three types theorem).
Let $G\curvearrowright B$ be such that $G$ is not a Boolean group acting freely on $B$. Let $s$ be the (possibly infinite) number of orbits of $G$ on $B$. Let $f$ be an almost minimal operation above $\overline{\langle G\rangle}$. Then, $f$ is of one of the following types:
  1. 1.
    a unary operation;
  2. 2.
    a binary operation;
  3. 3.
    a $k$-ary orbit/semiprojection for $3\leq k\leq s$.
Proof.
We need to consider the five possibilities for $f$ from Rosenberg’s five types theorem (Theorem 1.21). Lemma 2.3 and Lemma 2.2 tell us that $f$ cannot be a quasi-Malcev. Moreover, $f$ cannot be a quasimajority from Lemma 2.6. Finally, if $f$ is a quasi-semiprojection it must be an orbit/semiprojection from Lemma 2.1. Since $f$ is an orbit/semiprojection, its arity must be $\leq s$, where $s$ is the number of orbits of $G$. ∎
It is also easy to prove that essential $k$-ary orbit/semiprojections and various other operations introduced in Subsection 1.4 are always almost minimal by definition. We prove this in Lemma 2.10. From this observation, it also follows that there are almost minimal $k$-ary orbit/semiprojections for all $2\leq k\leq s$ (Proposition 2.11).
Definition 2.8.
Let $G\curvearrowright B$. A weak orbit/semiprojection is a $k$-ary operation $f$ for $k\geq 2$ such that for any distinct $i<j\leq k$, there is some $s(i,j)\leq k$ and $\alpha_{i,j}\in\overline{G}$ such that whenever $(a_{1},\dots,a_{k})$ is a $k$-tuple where $a_{i}$ and $a_{j}$ are in the same $G$-orbit,
$f(a_{1},\dots,a_{k})=\alpha_{i,j}(a_{s(i,j)}).$
Remark 2.9.
Let $G\curvearrowright B$. Suppose that $f$ is an orbit/semiprojection, a $G$-quasi-minority, an odd majority, or an odd Malcev. Then, $f$ is a weak orbit/semiprojection. This can be seen from the definitions of these operations.
Below we prove that essential weak orbit/semiprojections are almost minimal, thus showing that $G$-quasi-minorities, odd majorities, odd Malcev operations, and essential orbit/semiprojections are always almost minimal.
Lemma 2.10 (Collapse lemma).
Let $G\curvearrowright B$. Let $f:B^{k}\to B$ be an essential weak orbit/semiprojection such that $f(x,\dots,x)\in\overline{G}$. Then, $f$ is almost minimal above $\overline{\langle G\rangle}$.
Proof.
Let $r<k$. We show by induction on terms that all $r$-ary functions in $\langle G\cup\{f\}\rangle$ are essentially unary and arise from a function in $\overline{G}$ by adding dummy variables; taking the topological closure then yields the lemma. For the base case, the elements of $G$ and the projections satisfy the claim. In the induction step, it is sufficient to consider $f(g_{1},\dots,g_{k})$, where $g_{1},\dots,g_{k}$ are $r$-ary and satisfy our claim. Since all of the $g_{i}$ are essentially unary and $r<k$, two of them must depend on the same variable and so there are $1\leq i<j\leq k$ such that $g_{i}(x_{1},\dots,x_{r})$ and $g_{j}(x_{1},\dots,x_{r})$ always lie in the same $G$-orbit. In particular, since $f$ is a $k$-ary weak orbit/semiprojection, there is $\alpha_{i,j}\in\overline{G}$ and $s(i,j)\leq k$ such that
$f(g_{1},\dots,g_{k})(x_{1},\dots x_{r})=\alpha_{i,j}g_{s(i,j)}(x_{1},\dots x_{r}),$
Since by the inductive hypothesis, $g_{s(i,j)}$ is essentially unary and in $\langle\overline{G}\rangle$, so is $f(g_{1},\dots,g_{k})$. This completes the proof.
Proposition 2.11.
Let $G\curvearrowright B$, where $s$ is the number of orbits of $G$ on $B$. Then, for every $2\leq k\leq s$ there is a $k$-ary orbit/semiprojection which is almost minimal over $\overline{\langle G\rangle}$.
Proof.
We construct a $k$-ary quasi-semiprojection $f$ as follows. Let $(O_{r}|r<s)$ enumerate the $G$-orbits of $B$. Then, given $(a_{1},\dots,a_{k})\in B^{k}$, if any two of the $a_{i}$ are in the same orbit, $f(a_{1},\dots,a_{k})=a_{1}$. Otherwise, all of the $a_{i}$ belong to different orbits, and we let $f(a_{1},\dots,a_{k})=a_{j}$, where $a_{j}$ is the element appearing in the orbit with the largest index in our ordering. By construction, $f$ is an essential $k$-ary orbit/semiprojection. So almost minimality follows from Lemma 2.10. ∎
Lemma 2.12.
Let $G\curvearrowright B$ and suppose that $f$ is a $k$-ary orbit/semiprojection. Then, every function in $\overline{\langle G\cup\{f\}\rangle}\cap\mathcal{O}^{(k)}$ is a $k$-ary orbit/semiprojection.
Proof.
The proof has essentially the same structure as that of Lemma 2.10. It is sufficient to show the lemma holds for ${\langle G\cup\{f\}\rangle}\cap\mathcal{O}^{(k)}$. Suppose that $g_{1},\dots,g_{k}$ are $k$-ary orbit/semiprojections (we count among these also essentially unary operations). Now, suppose that $(a_{1},\dots,a_{k})$ is a tuple with two elements in the same orbit. We know that each of $g_{1},\dots,g_{k}$ depend on precisely one variable whenever this is the case. In particular, at least two of them will depend on the same variable, say $g_{i}$ and $g_{j}$ for $i<j$. Suppose that $f$ depends on the $l$th variable whenever two entries are in the same orbit and corresponds to $\alpha\in G$. Then,
$f(g_{1},\dots,g_{k})(a_{1},\dots,a_{k})=\alpha g_{l}(a_{1},\dots,a_{k}),$
and so $f(g_{1},\dots,g_{k})$ is also an orbit/semiprojection. In particular, we can prove that all operations in ${\langle G\cup\{f\}\rangle}\cap\mathcal{O}^{(k)}$ are orbit/semiprojections by induction on terms. ∎
3 Consequences of the three types theorem and oligomorphic clones
In this section, we explore some consequences of the three types Theorem 2.7 in an oligomorphic context, and more generally for clones whose unary operations correspond to $\overline{G}$ for $G\curvearrowright B$ non-trivial and not being the free action of a Boolean group on $B$.
3.1 Few reminders on oligomorphic permutation groups
Recall that for $B$ countable, we say that $G\curvearrowright B$ is oligomorphic when $G$ has finitely many orbits on $n$-tuples of $B$ for each $n\in\mathbb{N}$. By a theorem of Ryll-Nardzewski, Svenonius, and Engeler, oligomorphic permutation groups correspond to the automorphism groups of countable $\omega$-categorical structures, where a structure $B$ is $\omega$-categorical if its first-order theory has a unique countable model up to isomorphism [fathodges, Theorem 7.3.1].
Many natural examples of $\omega$-categorical structures come from countable homogeneous structures: relational structures where any isomorphism between finite substructures extends to an automorphism of the whole structure. These in turn correspond to well-behaved classes of finite structures (Fraïssé classes): by Fraïssé’s Theorem for any Fraïssé class $\mathcal{F}$ there is a countable homogeneous structure whose age, i.e. class of finite substructures is $\mathcal{F}$, and vice versa, the age of any countable homogeneous structure is a Fraïssé class. We call the homogeneous structure associated to the Fraïssé class $\mathcal{F}$ the Fraïssé limit of $\mathcal{F}$. Examples of Fraïssé classes are the finite linear orders, finite graphs, and finite triangle-free graphs, which have as associated Fraïssé limits, $(\mathbb{Q},<)$, the random graph, and the generic triangle-free graph. Homogeneous structures in a finite language are always $\omega$-categorical. We refer the reader to [homogeneous] for a survey on homogeneous structures.
Work on infinite domain constraint satisfaction is especially focused on homogeneous structures whose Fraïssé class is defined by omitting embeddings of finitely many structures in a given finite relational signature. These are called finitely bounded homogeneous structures. A conjecture of Bodirsky and Pinsker asks whether a similar complexity dichotomy to the Bulatov [bulatov2017dichotomy] and Zhuk [zhuk2020proof] holds for first-order reducts of finitely bounded homogeneous structures. We refer the reader to Bodirsky’s book [BodCSP] and to a review of the second author [infinitesheep] for more on infinite domain constraint satisfaction problems.
An important feature of $\omega$-categorical structures in a finite relational language is that if their polymorphism clone contains an essential operation, then it will also contain a minimal operation above $\overline{\langle\mathrm{Aut}(B)\rangle}$.
Fact 3.1 (Existence of minimal clones, oligomorphic case, [BodChen]).
Let $\mathcal{L}$ be a finite relational language. Let $B$ be a countable $\omega$-categorical $\mathcal{L}$-structure. Let $\mathcal{D}$ be a closed clone properly containing $\overline{\langle\mathrm{Aut}(B)\rangle}$. Then, $\mathcal{D}$ contains a closed clone $\mathcal{C}$ which is minimal above $\overline{\langle\mathrm{Aut}(B)\rangle}$.
3.2 Essential polymorphisms in $\omega$-categorical structures
In this subsection, we look at Question 1 of Bodirsky, which asks whether an $\omega$-categorical model-complete core with an essential polymorphism also has a binary essential polymorphism. Whilst we give a negative answer to it in Corollary 3.8, we prove in Corollary 3.7 that the answer is positive for $\omega$-categorical structures with less than two orbits. In Subsection 3.3, we find weaker assumptions that imply the existence of binary essential polymorphisms. The results in this subsection are direct consequences of Theorem 2.7.
Definition 3.2.
Let $G\curvearrowright B$ and $\mathcal{C}$ be a clone strictly containing $\overline{\langle G\rangle}$. We say that $\mathcal{C}$ is a core clone with respect to $G$ if $\mathcal{C}\cap\mathcal{O}^{(1)}=\overline{G}$.
Definition 3.3.
We say that a countable $\omega$-categorical structure $B$ is a model-complete core if $\mathrm{Pol}(B)$ is a core clone.
Remark 3.4.
Let $B$ be a countable $\omega$-categorical structure with automorphism group $G$. Any closed subclone of $\mathrm{Pol}(B)$ containing $\overline{\langle G\rangle}$ is the polymorphism clone of some first-order reduct of $B$ (essentially by Fact 1.7 and Ryll-Nardzewski’s Theorem [fathodges], c.f. Theorem 6.1.13 [BodCSP]). In particular, if $\mathcal{C}=\overline{\langle G\cup\{f\}\rangle}$ for $f$ almost minimal above $\overline{\langle G\rangle}$, then $\mathcal{C}$ is the polymorphism clone of a model-complete core which is a first-order reduct of $B$.
Remark 3.5.
No oligomorphic permutation group $G\curvearrowright B$ is a Boolean group acting freely on a set. So, Theorem 2.7 applies to this context. To see this, note that the action of an oligomorphic permutation group is never free. Take a point $a\in B$. The action of the stabilizer of $a$, $G_{a}$ on $B$ still has finitely many orbits by oligomorphicity. Hence, there are two points $b,c\in B$ such that $ab$ and $ac$ lie in the same $G$-orbit. In particular, this means that there is a non-trivial automorphism fixing $a$, contradicting freeness.
Corollary 3.6.
Let $G\curvearrowright B$ not be the free action of a Boolean group and have $\leq 2$ orbits. Let $\mathcal{C}$ be a core clone with respect to $G$. Then, $\mathcal{C}$ contains a binary essential polymorphism.
Proof.
By Lemma 1.20, since $\mathcal{C}\supsetneq\overline{\langle G\rangle}$, $\mathcal{C}$ has an almost minimal operation $f$. Since $\mathcal{C}\cap\mathcal{O}^{(1)}=\langle G\rangle$, $f$ cannot be unary. By the three types theorem (Theorem 2.7), since $s\leq 2$, $f$ has to be binary. Clearly, $f$ is essential. ∎
Corollary 3.7.
Let $B$ be an $\omega$-categorical countable model-complete core such that $\mathrm{Aut}(B)$ has $\leq 2$ orbits. Then, if $\mathrm{Pol}(B)$ has an essential polymorphism, it also has a binary essential polymorphism.
Proof.
Follows from Corollary 3.6 and the definition of an $\omega$-categorical model-complete core. ∎
Corollary 3.8.
Let $B$ be a countable $\omega$-categorical structure such that $\mathrm{Aut}(B)$ has $s\geq 3$ orbits on $B$. Then, for each $3\leq k\leq s,B$ has a first-order reduct which is a model-complete core and such that it has a $k$-ary essential polymorphism and no essential polymorphisms of lower arity.
Proof.
By Theorem 2.11, for each $3\leq k\leq s$, there is a closed clone $\mathcal{C}$ above $\overline{\langle\mathrm{Aut}(B)\rangle}$ generated by an $k$-orbit/semiprojection and $\overline{\langle\mathrm{Aut}(B)\rangle}$. By Remark 3.4, $\mathcal{C}$ is the polymorphism clone of a first-order reduct of $B$ which is a model-complete core. ∎
Corollary 3.8 gives a negative answer to Question 1. However, Corollary 3.7 yields that Question 1 has a positive answer whenever $B$ is an $\omega$-categorical model-complete core whose automorphism group has less than two orbits. In Section 3.3, in particular in Theorem 3.14, we see that for $B$ an $\omega$-categorical model-complete core whose CSP is not NP-hard we can always find a binary essential operation in $\mathrm{Pol}(B)$.
3.3 Why easy problems lie above binary minimal polymorphisms
In this subsection, we show Theorem 3.14, which says that as long as $\mathcal{C}$ is a core clone above a suitable permutation group $G\curvearrowright B$ without a continuous $h1$-homomorphism to the clone of projections it will have a binary essential polymorphism. This is especially helpful for the study of $\mathrm{CSP}$s of $\omega$-categorical structures since such a homomorphism is known to imply hardness of the CSP [Wonderland]. This is an interesting disanalogy with the case of idempotent clones (i.e. core clones above the trivial group), where on any domain there are clones with no essential binary operations but no $h1$-homomorphism to projections (e.g. any clone generated by a majority or minority operation).
Notation 3.9.
By $\mathcal{P}_{\{0,1\}}$ we denote the clone of projections on a two-element domain.
Definition 3.10.
A height $\bm{1}$-identity, shortened "$h1$-identity", is an identity with exactly one function symbol on each side.
The following is an example of an $h1$-identity:
$f(x,x,y)\approx g(x,y).$
The following are not $h1$-identities:
$f(x,y,f(x,y,z))\approx g(x,y),\quad g(x,\beta y)\approx g(x,\alpha y),\quad f(x,x,x)\approx x.$
The first two identities have two function symbols on one of the two sides of the equation, and the third has no function symbol on one of the two sides. Hence, none of them are $h1$-identities.
Definition 3.11.
A $\bm{h1}$-homomorphism (also known as a minion homomorphism) between function clones $\eta:\mathcal{C}\to\mathcal{D}$ is a map preserving $h1$-identities. Equivalently, it is a map preserving arities and compositions with projections.
From [Wonderland, Theorem 1.8], we know that if $\mathcal{C}$ is the polymorphism clone of a finite or $\omega$-categorical structure $B$ and it has a uniformly continuous $h1$-homomorphism to $\mathcal{P}_{\{0,1\}}$, then $\mathrm{CSP}(B)$ is NP-hard. Here, uniform continuity is in terms of the pointwise convergence topology. In particular, for $\mathcal{C}$ a clone on $B$, a map $\eta:\mathcal{C}\to\mathcal{D}$ preserving arities is uniformly continuous if there is some finite $B^{\prime}\subseteq B$ such that $f_{\upharpoonright B^{\prime}}=g_{\upharpoonright B^{\prime}}$ implies $\eta(f)=\eta(g)$.
Definition 3.12.
Let $\mathcal{C}$ be a clone. Let $\xi:\mathcal{C}\cap\mathcal{O}^{\leq(3)}\to\mathcal{P}_{\{0,1\}}$ be a map preserving arities. The standard extension of $\xi$ is the map $\xi^{\prime}:\mathcal{C}\to\mathcal{P}_{\{0,1\}}$ defined as follows:
Let $f\in\mathcal{C}$ be an $n$-ary operation. Let $(a_{1},\dots,a_{n}):=a\in\{0,1\}^{n}$. Write $f_{a}(x,y)$ for the binary operation induced by $f$ substituting the variable $x_{i}$ with $x$ if $a_{i}=0$ and with $y$ otherwise. We then, define
$\xi^{\prime}(f)(a):=\xi(f_{a})(0,1).$
From [irrelevant, Proposition 6.8] we know the following:
Fact 3.13.
Let $\mathcal{C}$ be a clone. Suppose that $\xi:\mathcal{C}\cap\mathcal{O}^{\leq(3)}\to\mathcal{P}_{\{0,1\}}$ is a map preserving arities and $h1$ identities. Then, the standard extension $\xi^{\prime}:\mathcal{C}\to\mathcal{P}_{\{0,1\}}$ is an $h1$-homomorphism.
In particular, Fact 3.13 tells us that satisfying any non-trivial $h1$-identity implies having some essential operation of arity $\leq 3$. We prove that even extremely small amounts of symmetry force the existence of a binary essential operation:
Theorem 3.14.
Let $G\curvearrowright B$ be such that $G$ is not a Boolean group acting freely on $B$. Suppose that $\mathcal{C}$ is a core clone with respect to $G$. Suppose that $\mathcal{C}$ has no uniformly continuous h1-homomorphism to the clone of projections on a two-element set $\mathcal{P}_{\{0,1\}}$. Then, $\mathcal{C}$ contains a binary essential operation almost minimal above $\overline{\langle G\rangle}$.
Proof.
Note that if $f\in\mathcal{C}\cap\mathcal{O}^{(2)}\supsetneq\overline{\langle G\rangle}\cap\mathcal{O}^{(2)}$, since $\mathcal{C}$ is a core clone with respect to $G$, the operation $f$ must be binary and almost minimal. Hence, suppose that $\mathcal{C}\cap\mathcal{O}^{(2)}=\overline{\langle G\rangle}\cap\mathcal{O}^{(2)}$. Then, all ternary operations in $\mathcal{C}\setminus\overline{\langle G\rangle}$ must be almost minimal. In particular, $\mathcal{C}\cap\mathcal{O}^{(3)}$ consists entirely of essentially unary operations and orbit/semiprojections.
Hence, we can construct a map $\xi:\mathcal{C}\cap\mathcal{O}^{\leq(3)}\to\mathcal{P}_{\{0,1\}}$ preserving arities and $h1$-identities. Firstly, $\xi$ sends the elements of $\overline{G}$ to the identity. Essentially unary operations depending on their $i$th variable are sent to the projection of the same arity to the $i$th coordinate. The (potential) ternary orbit/semiprojections, which when two elements are in the same orbit depend on the $i$th variable are sent by $\xi$ to the projection to the $i$th variable of the same arity. By construction $\xi$ preserves aritites. It is easy to see that $\xi$ preserves compositions with projections and so all $h1$ identities. Hence, $\xi$ is an $h1$-homomorphism to $\mathcal{P}_{\{0,1\}}$.
By Fact 3.13, the standard extension $\xi^{\prime}:\mathcal{C}\to\mathcal{P}_{\{0,1\}}$ is an $h1$-homomorphism. To prove uniform continuity, we need to show there is some finite set $B^{\prime}\subset B$ such that for all $f,g\in\mathcal{C}$, if $f$ and $g$ agree on $B^{\prime}$, then they are sent to the same operation in $\mathcal{P}_{\{0,1\}}$ by $\xi^{\prime}$. By assumption $\mathcal{C}\cap\mathcal{O}^{(2)}=\overline{\langle G\rangle}\cap\mathcal{O}^{(2)}$ and so all binary functions in $\mathcal{C}$ are essentially unary. Hence, considering the construction of $\xi^{\prime}$ and the definition of $\xi$, we can see that for two functions $f,g\in\mathcal{C}$ of the same arity, $\xi^{\prime}(f)=\xi^{\prime}(g)$ if and only if for each $a\in\{0,1\}^{n},f_{a}(x,y)$ and $g_{a}(x,y)$ depend on the same variable. One can determine for all $a\in\{0,1\}^{n}$ whether $f_{a}(x,y)$ and $g_{a}(x,y)$ depend on the same variable by only studying the behaviour of $f$ and $g$ on tuples consisting of two distinct elements $b,c\in B$. In fact, being essentially unary, $f_{a}(b,c)$ will be equal to either $f_{\overline{0}}(b,c)$ or $f_{\overline{1}}(b,c)$, where $\overline{0}=(0,\dots,0),\overline{1}=(1,\dots,1)\in\{0,1\}^{n}$. Since $\mathcal{C}\cap\mathcal{O}^{(1)}$ consists of injective functions, $f_{a}(b,c)$ cannot be equal to both. Hence, whether $f_{a}(b,c)$ equals $f_{\overline{0}}(b,c)$ or $f_{\overline{1}}(b,c)$ determines entirely on which variable $f_{a}$ depends. Thus, if $f$ and $g$ agree on all tuples from $\{b,c\}$, $\xi^{\prime}(f)=\xi^{\prime}(g)$, yielding uniform continuity and concluding the argument. ∎
Remark 3.15.
As pointed out earlier, over any domain there is an idempotent clone $\mathcal{C}$ with no $h1$-homomorphism to $\mathcal{P}_{\{0,1\}}$. If $G\curvearrowright B$ is a Boolean group acting freely on $B$, from Proposition 4.8 there is always an almost minimal $G$-quasi-minority $m$ above $\langle G\rangle$. This will not have a $h1$-homomorphism to projections since a minority satisfies a set of non-trivial $h1$-identities which cannot be satisfied by a projection. But by almost minimality $\overline{\langle G\cup\{m\}\rangle}$ will not have any binary essential operations.
Since a uniformly continuous $h1$-homomorphism to projections implies NP-hardness in a finite or $\omega$-categorical setting [Wonderland, Theorem 1.8], we can deduce that if $\mathrm{CSP}(B)$ is not NP-hard, there must be some binary essential polymorphism $f\in\mathrm{Pol}(B)$ witnessing this. In particular,
Corollary 3.16.
Suppose that $B$ is a finite or countable $\omega$-categorical model-complete core and $\mathrm{Aut}(B)\curvearrowright B$ is not the free action of a Boolean group on $B$ (this is always the case if $B$ is $\omega$-categorical). Suppose that $\mathrm{CSP}(B)$ is not NP-hard. Then, $\mathrm{Pol}(B)$ contains a binary essential polymorphism.
Obviously, if P=NP, the above statement is vacuously true.
3.4 The orbital extension property and its failure
Finding binary essential operations given some essential operation in $\mathrm{Pol}(B)$ for $B$ $\omega$-categorical is useful for several arguments in the context of infinite-domain CSPs [BodKara, Smoothapp, bodirsky2020complexity]. For this reason, previous authors already developed techniques to find binary essential operations given an essential operation [BodKara, minrandom, Smoothapp]. The standard technique made use of the orbital extension property:
Definition 3.17.
We say that $G\curvearrowright B$ has the orbital extension property if there is an orbital (i.e. orbit of $G$ on $B^{2}$) $O$ such that for any $u,v\in B$ there is $z\in B$ such that $(u,z),(v,z)\in O$.
Note that having the orbital extension property implies transitivity. A lot of transitive $\omega$-categorical structures have the orbital extension property. For example, the following classes of structures have the orbital extension property:
  • any homogeneous structure with free amalgamation;
  • any $2$-transitive $\omega$-categorical structure;
  • any $\omega$-categorical structure $M$ with an $\mathrm{Aut}(M)$-invariant $1$-type over $M$;
  • any homogeneous transitive structure $B$ where the set of minimal bounds for $\mathrm{Age}(B)$ has no structures of size $3$.
An example of a transitive $\omega$-categorical structure which does not have the orbital extension property is $K_{\omega,\omega}$, the complete bipartite graph where both parts of the partition are countable.
As mentioned, one can prove that if $A$ is a first-order reduct of an $\omega$-categorical structure $B$ such that $\mathrm{Aut}(B)$ has the orbital extension property, if $\mathrm{Pol}(A)$ is not essentially unary, it must contain a binary essential operation. The following was first observed in a weaker form in [BodKara, Lemma 5], and then strengthened in [minrandom, Lemma 40] and [Smoothapp, Proposition 22]:
Fact 3.18.
Let $\mathcal{C}$ be a closed clone with an essential operation and containing a permutation group $G\curvearrowright B$ with the orbital extension property. Then, $\mathcal{C}$ contains a binary essential operation.
As noted above, a lot of transitive $\omega$-categorical structures have the orbital extension property. Indeed, it was not clear whether one could prove that all primitive $\omega$-categorical structures have the orbital extension property. For this reason, in his book [BodCSP], Bodirsky asks the following question mentioned in the introduction:
Question 2.
(Question 14.2.1 (2) in [BodCSP]) Does every primitive oligomorphic permutation group have the orbital extension property?
We shall see in Proposition 3.22 that the answer to Question 2 is negative. This yields that our results in Corollary 3.7 offer a strict improvement with respect to previous techniques. On one hand, our results made Fact 3.18 redundant and cover non-primitive and non-transitive structures: now we know that as long as $B$ is an $\omega$-categorical core with $\mathrm{Aut}(B)$ having less than two orbits, if $\mathrm{Pol}(B)$ is not essentially unary it must contain a binary essential operation. Indeed Theorem 3.14 even tells us that this will happen as long as $\mathrm{Pol}(B)$ does not have a uniformly continous $h1$-homomorphism to projections, which for the purposes of studying CSPs is everything one needs. On the other hand, we will now prove that the orbital extension property fails even for some very natural binary finitely bounded primitive homogeneous structures, showing that techniques using this property cannot be generalised to all primitive $\omega$-categorical model-complete cores, unlike our results.
The structures which will yield a negative answer to Question 2 are homogeneous $2$/multitournaments:
Definition 3.19.
A $\bm{2}$-multitournament is a relational structure $\mathcal{M}$ in a language with two binary relations {$\right\}$ such that for any two points $a,b\in M,$ exactly one of
[…]
holds. We can think of a $2$-multitournament as a tournament where every arc is coloured of one of two colours.
The family of primitive $3$-constrained homogeneous $2$-multitournaments has recently been classified by Cherlin [Cherlinclass, Table 18.1]. We shall see that two such multitournaments do not have the orbital extension property. In order to introduce them we need to set up some notation to denote which substructures we are omitting from the Fraïssé class of $2$-multitournaments yielding our desired counterexamples.
Notation 3.20.
For $i,j,k\in\{1,2\}$, and $C(i,j,k)$ denotes the $2$-multitournament of size $3$ consisting of an oriented $3$-cycle given by three points $a,b,c$ such that
[…]
Similarly, $L(i,j,k)$ denotes the $2$-multitournament of size $3$ where the three vertices $a,b,c$ are such that
[…]
Definition 3.21.
The homogeneous $2$-multitournaments $\widetilde{\mathbb{S}(3)}$ and $\mathbb{S}(4)$ are given by the Fraïssé limit of the classes of $2$/multitournaments omitting embeddings of the $2$/multitournaments on three vertices described by Table 1.
      Structure       omitted instances of $C(i,j,k)$       omitted instances of $L(i,j,k)$
      $\widetilde{\mathbb{S}(3)}$       111, 222       111, 122, 212
      $\mathbb{S}(4)$       111, 112       121, 211, 221, 222
Table 1: In the row of a given structure $S$ we list the $2$-multitournaments of size $3$ such that the class of $2$-multitournaments omitting them yields $S$ as a Fraïssé limit. By having "111" written in the "omitted instances of $C(i,j,k)$" column on the $\widetilde{\mathbb{S}(3)}$-row, we mean that the $2$-multitournament of size $3$ given by $C(1,1,1)$ is omitted from the class of finite $2$-multitournaments yielding $\widetilde{\mathbb{S}(3)}$ as a Fraïssé limit. In particular, in Figure 2 we represent the omitted triangles in $\widetilde{\mathbb{S}(3)}$ of the form $L(i,j,k)$.
[…]
Figure 2: pictorial representation of the $L(i,j,k)$-omitted triangles in $\widetilde{\mathbb{S}(3)}$.
Proposition 3.22.
The automorphism groups of the homogeneous $2$-multitournaments $\widetilde{\mathbb{S}(3)}$ and $\mathbb{S}(4)$ do not have the orbital extension property.
Proof.
We run the proof for $\widetilde{\mathbb{S}(3)}$ since the proof for $\mathbb{S}(4)$ is virtually identical. There are four possible orbits of pairs for which the orbital extension property could hold. The types of $\widetilde{\mathbb{S}(3)}$ corresponding to these orbitals are isolated by the following formulas:
[…]
As pointed out earlier, these are primitive binary finitely bounded homogeneous structures [Cherlinclass], thus yielding a counterexample to Question 2.
4 Almost minimal operations above a Boolean group
In this section we classify almost minimal operations above a Boolean group $G\curvearrowright B$ of size $>2$ in Theorem 4.7. Compared to the three types Theorem 2.7, we have an extra kind of almost minimal operation: $G$-quasi-minorities (Definition 1.24). The case of $|G|=2$ will be treated later in Section 6.
Remark 4.1.
Note that if $G\curvearrowright B$ freely, then $\overline{G}=G$ in the pointwise convergence topology on $\mathcal{O}^{(1)}$. This is just because by freeness of the action for $\alpha,\beta\in G,a\in B,\alpha a=\beta a$ implies that $\alpha=\beta$, so the only convergent sequences from $G$ are those which are eventually constant. Moreover, one can see by induction on terms that $\langle G\rangle=\overline{\langle G\rangle}$.
Lemma 4.2.
Let $G\curvearrowright B$ where $|G|>2$. Then, there is no quasi-majority almost minimal above $\overline{\langle G\rangle}$.
Proof.
Suppose by contradiction that $m$ is a quasi-majority almost minimal above $\overline{\langle G\rangle}$. Since $|G|>2$, take $\alpha,\beta\in G\setminus\{1\}$ distinct. Lemma 2.4 implies $m(x,\gamma x,z)\approx m(z,z,z)\approx m(z,\gamma x,x)$ for any $\gamma\in G\setminus\{1\}$. Let $u$ be a new variable. Setting $x=u,\gamma=\alpha,z=\beta u$ in the first equation yields $m(u,\alpha u,\beta u)\approx m(\beta u,\beta u,\beta u)$; on the other hand, setting $x=\beta u,z=u,\gamma=\alpha\circ\beta^{-1}$ in the second equation yields $m(u,\alpha u,\beta u)\approx m(u,u,u)$, contradicting injectivity of the map $x\mapsto m(x,x,x)$. ∎
Lemma 4.3.
Let $G\curvearrowright B$ where $G$ is a Boolean group. Let $\alpha\in G\setminus\{1\}$. Suppose that $M(x,y,z)$ is quasi-Malcev and almost minimal above ${\langle G\rangle}$. Then, $h(x,y,z):=M(x,\alpha y,z)$ is not a quasi-semiprojection.
Proof.
The operation $h$ is also almost minimal above ${\langle G\rangle}$. Suppose by contradiction $h(x,y,z)$ is a quasi-semiprojection. Suppose that when two variables are the same $h$ depends on the first coordinate and equals the unary function $g\in{\langle G\rangle}$. Then,
$g(y)\approx h(y,\alpha y,y)\approx M(y,y,y)\approx M(\alpha y,\alpha y,y)\approx h(\alpha y,y,y)\approx g(\alpha y)$
By injectivity of $g$, this cannot happen. The same argument works for showing that $h$ cannot depend on the third coordinate. So, suppose that $h(x,y,z)$ depends on the second coordinate when two variables are identified. Using the fact that $G$ is Boolean, we get that
$g(y)\approx h(\alpha y,y,\alpha y)\approx M(\alpha y,\alpha y,\alpha y)\approx M(\alpha y,y,y)\approx h(\alpha y,\alpha y,y)\approx g(\alpha y),$
again contradicting injectivity of $g\in{\langle G\rangle}$. Hence, $h$ cannot be a quasi-semiprojection.
Lemma 4.4.
Let $G\curvearrowright B$ and let $\alpha\in G\setminus\{1\}$. Suppose that $M(x,y,z)$ is quasi-Malcev and almost minimal above ${\langle G\rangle}$. If some permutation of the variables of $h(x,y,z)=M(x,\alpha y,z)$ is quasi-Malcev, then so is $h$ itself.
Proof.
Suppose $h$ is not quasi-Malcev itself, but some function obtained by permuting its variables is. Then $h(x,y,x)\approx h(y,y,y)$, and hence
$h(y,y,y)\approx h(\alpha y,y,\alpha y)\approx M(\alpha y,\alpha y,\alpha y)\approx M(\alpha y,y,y)\approx h(\alpha y,\alpha y,y)\;.$
By almost minimality, the function $h(x,x,y)$ is essentially unary, and depends on one of its variables injectively. By the above, this variable has to be $y$, and we moreover see $h(x,x,y)\approx h(y,y,y)$. The same argument shows that $h(y,x,x)\approx h(y,y,y)$, so $h$ is quasi-Malcev, a contradiction.
Recall the notion of a $G$-quasi-minority from Definition 1.24.
Lemma 4.5.
Let $G\curvearrowright B$ where $G$ is a Boolean group. Let $M(x,y,z)$ be a quasi-Malcev operation almost minimal above ${\langle G\rangle}$. Let $\beta\in G\setminus\{1\}$, and suppose that $M(x,\beta y,z)$ is quasi-Malcev. Then,
$M(x,\beta x,y)$ $\approx M(\beta y,\beta y,\beta y)\approx M(y,\beta x,x)$ (7)
$M(x,y,\beta x)$ $\approx M(\beta y,\beta y,\beta y)$ (8)
In particular, $M(x,y,z)$ is a $G$-quasi-minority.
Proof.
By assumption, the function $h(x,y,z)=M(x,\beta y,z)$ is quasi-Malcev. We thus obtain the first equation in 7:
$M(x,\beta x,y)\approx h(x,x,y)\approx h(y,y,y)\approx h(y,\beta y,\beta y)\approx M(y,y,\beta y)\approx M(\beta y,\beta y,\beta y).$
By symmetry, the same argument also yields the second equation in 7.
For equation 8, note that by almost minimality of $M(x,y,z),\ l(x,y)=M(x,y,\beta x)$ is in ${\langle G\rangle}$ and so it can only depend on either the first or the second coordinate. We have
$l(x,x)\approx M(x,x,\beta x)\approx\beta x\qquad\text{and}\qquad l(x,\beta x)\approx M(x,\beta x,\beta x)\approx x\;,$
hence that coordinate is the second one. This implies identity 8 since
$M(x,y,\beta x)\approx l(x,y)\approx l(y,y)\approx M(y,y,\beta y)\approx M(\beta y,\beta y,\beta y)\;.$
In particular, we can conclude that $M$ is a $G$-quasi-minority since it satisfies equation 1. ∎
Corollary 4.6.
Let $G\curvearrowright B$ where $G$ is a Boolean group with $|G|>2$. Let $M(x,y,z)$ be a quasi-Malcev operation almost minimal above ${\langle G\rangle}$. Then, $M(x,y,z)$ is a $G$-quasi-minority.
Proof.
Pick an arbitrary $\beta\in G\setminus\{1\}$. The function $M(x,\beta y,z)$ is almost minimal above $\langle G\rangle$. By Lemma 4.2 and since $|G|>2$, this function cannot be a quasi-majority, and by Lemma 4.3, it cannot be a quasi-semiprojection. Hence, some permutation of its variables is quasi-Malcev, and so it is quasi-Malcev itself by Lemma 4.4. Thus, Lemma 4.5 applies, and we see that $M(x,y,z)$ is a $G$-quasi-minority.
Theorem 4.7 (Boolean case).
Let $G\curvearrowright B$ be a Boolean group acting freely on $B$ with $s$-many orbits (where $s$ is possibly infinite) and $|G|>2$. Let $f$ be an almost minimal operation above $\langle G\rangle$. Then, $f$ is of one of the following types:
  1. 1.
    a unary operation;
  2. 2.
    a binary operation;
  3. 3.
    a ternary $G$-quasi-minority;
  4. 4.
    a $k$-ary orbit/semiprojection for $3\leq k\leq s$.
Proof.
It suffices to examine $f$ in the light of Rosenberg’s five types theorem 1.21 and apply our additional results for Boolean groups of order greater than two. Lemma 4.2 tells us that $f$ cannot be a quasi-majority. Corollary 4.6 tells us that if $f$ is quasi-Malcev, then it must be a $G$-quasi-minority. ∎
Proposition 4.8.
Let $G\curvearrowright B$ where $G$ is a Boolean group acting freely on $B$. Then there exists a $G$-quasi-minority almost minimal above $\langle G\rangle$.
Proof.
For any triple $a=(a_{1},a_{2},a_{3})$ of elements which are not in the same orbit, we consider $(i,j)$ smallest lexicographically such that $a_{i},a_{j}$ are in the same orbit, and let $k$ be the remaining coordinate. Since the action of $G$ is free, there exists a unique $\alpha\in G$ such that $a_{j}=\alpha a_{i}$; we set $f(a_{1},a_{2},a_{3})=\alpha a_{k}$ and extend $f$ to $B^{3}$ arbitrarily. Note that by definition, $f(x,x,x)\approx x$. To see that $f$ is a $G$-quasi-minority, consider first a triple $(a_{1},a_{2},a_{3})$ such that precisely two of its elements $a_{i},a_{j}$ belong to the same orbit, i.e., $a_{i}=\alpha a_{j}$ for a unique $\alpha\in G$. Then also $a_{j}=\alpha a_{i}$ since $G$ is Boolean. By definition, for the remaining coordinate $k\notin\{i,j\}$ we have $f(a_{1},a_{2},a_{3})=\alpha a_{k}$ independently of whether $i<j$. Now consider a triple all three of whose elements belong to the same orbit, i.e., $(a_{1},a_{2},a_{3})=(\gamma_{1}a,\gamma_{2}a,\gamma_{3}a)$ for some $\gamma_{i}\in G$ and $a\in B$. Then, using that $G$ is Boolean, we have $f(a_{1},a_{2},a_{3})=\gamma_{1}\gamma_{2}a_{3}$ by definition. This equals $\gamma_{1}\gamma_{2}\gamma_{3}a$, and since $G$ is Boolean, in turn also $\gamma_{1}\gamma_{3}a_{2}$ and $\gamma_{2}\gamma_{3}a_{1}$, showing that $f$ is indeed a $G$-quasi-minority. Almost minimality follows from Remark 2.9 and Lemma 2.10. ∎
5 Minimal $G$-quasi-minorities
In this section, we characterise minimal $G$-quasi-minorities as $G$-invariant Boolean Steiner $3$-quasigroups (Definition 1.28) in Theorem 5.21. Subsection 5.1, is dedicated to proving this whilst in Subsection 5.2 we further characterise the behaviour of these operations to obtain necessary and sufficient conditions for their existence (Corollary 5.32) as well as an explicit understanding of their behaviour (Definitions 5.23 and 5.27). As an intermediate step towards understanding minimal $G$-quasi-minorities we also classify strictly almost minimal $G$-quasi-minorities 5.8, i.e. $G$-quasi-minorities which only generate ternary operations that generate them back. These can also be nicely characterised as $G$-invariant Steiner $3$-quasigroup (Definition 5.5).
5.1 $G$-invariant Steiner $3$-quasigroups
Definition 5.1.
A Steiner $3$-quasigroup is an idempotent symmetric minority $m$ satisfying the equation
$m(x,y,m(x,y,z))\approx z.$
Definition 5.2.
A Steiner quadruple system, shortened to SQS, is a uniform $4$-hypergraph such that any three vertices are contained in a unique $4$-hyperedge.
Steiner $3$-quasigroups on a given domain $B$ naturally correspond to Steiner quadruple system on $B$: the $4$-ary relation
$(m(x_{1},x_{2},x_{3})=x_{4})\wedge\bigwedge_{i<j\leq 4}x_{i}\neq x_{j}$
yields a Steiner quadruple system on $B$ since one can prove by SQS this is a symmetric $4$-ary relation such that every three vertices are in a unique $4$-hyperedge (we essentially prove this again in Lemma 5.10). The converse procedure yields a Steiner $3$-quasigroup for any SQS on $B$. The connection between Steiner quadruple systems and Steiner $3$-quasigroups is heavily studied in universal algebra and design theory, starting with the work of [quackenbush1975algebraic, armanious1980algebraische]. We refer the reader to [quackenbush1999nilpotent] and [lindner1978steiner, Section 4] for discussions of this connection. In earlier literature these operations are frequently called ’Steiner skeins’.
An important identity that a Steiner $3$-quasigroup might satisfy is the Boolean law:
$m(x,y,m(z,y,w))\approx m(x,z,w).$ (Bool)
The Boolean law says that in the Steiner quadruple system associated with $m$, whenever two $4$-hyperedges $H_{1}$ and $H_{2}$ intersect in two vertices, the remaining four vertices in $H_{1}\cup H_{2}$ form another $4$-hyperedge.
Definition 5.3.
A Steiner $3$-quasigroup satisfying the Boolean law (Bool) is called Boolean.
Remark 5.4.
It is easy to see that given a Boolean Steiner $3$-quasigroup, fixing any vertex $e\in B$, the structure $(B,+,e)$ with operation $a+b:=m(a,b,e)$ is a Boolean group and for all $a,b,c\in B$,
$m(a,b,c)=a+b+c$
with respect to this operation [quackenbush1999nilpotent]. The associated SQSs are precisely Kirkman’s Steiner quadruple systems [kirkman1847problem].
Definition 5.5.
Let $G\curvearrowright B$ be a Boolean group acting freely on $B$. A $\bm{G}$-invariant Steiner $\bm{3}$-quasigroup is a Steiner $3$-quasigroup satisfying the $\bm{G}$-invariance identity: for all $\alpha,\beta,\gamma\in G$
$m(\alpha x,\beta y,\gamma z)\approx\alpha\beta\gamma m(x,y,z).$ (Inv)
Lemma 5.6.
Let $G\curvearrowright B$ be a Boolean group acting freely on $B$. Suppose that $\mathfrak{m}$ is a $G$-quasi-minority. Let $\alpha,\beta,\gamma\in G$. The operation $\mathfrak{m}^{\prime}$ given by
$\mathfrak{m}^{\prime}(x,y,z):=\mathfrak{m}(x,y,\mathfrak{m}(\alpha x,\beta y,\gamma z))$
is an orbit/semiprojection depending on the last coordinate.
Proof.
Say that $\mathfrak{m}(x,x,x)\approx\eta x$ for $\eta\in G$. Since $\mathfrak{m}$ is almost minimal, so must be $\mathfrak{m}^{\prime}$. For $\delta\in G$, we have the following:
$\mathfrak{m}^{\prime}(x,y,\delta x)$ $:=\mathfrak{m}(x,y,\mathfrak{m}(\alpha x,\beta y,\delta\gamma x))\approx\mathfrak{m}(x,y,\eta\alpha\beta\gamma\delta y)\approx\alpha\beta\gamma\delta x;$ (9)
$\mathfrak{m}^{\prime}(x,\delta x,y)$ $:=\mathfrak{m}(x,\delta x,\mathfrak{m}(\alpha x,\beta\delta x,\gamma y))\approx\mathfrak{m}(x,\delta x,\eta\alpha\beta\gamma\delta y)\approx\alpha\beta\gamma y;$ (10)
$\mathfrak{m}^{\prime}(y,x,\delta x)$ $:=\mathfrak{m}(y,x,\mathfrak{m}(\alpha y,\beta x,\delta\gamma x))\approx\mathfrak{m}(y,x,\eta\alpha\beta\gamma\delta y)\approx\alpha\beta\gamma\delta x.$ (11)
In particular, whenever two entries are in the same $G$-orbit, $\mathfrak{m}^{\prime}$ depends on the last coordinate. ∎
Remark 5.7.
The above Lemma implies that if $\mathfrak{m}$ is a minimal $G$-quasi-minority, then $\mathfrak{m}^{\prime}\in\langle G\rangle$. This is because from Lemma 2.12 an orbit/semiprojection only generates more orbit/semiprojections. Hence, for $\mathfrak{m}$ to be minimal, $\mathfrak{m}^{\prime}$ cannot be an essential orbit/semiprojection and must be essentially unary. This yields some very strong constraints on the behaviour of $\mathfrak{m}$ itself, which we will exploit in the rest of this section.
We begin by studying a notion which is strictly weaker than minimality and strictly stronger than almost minimality:
Definition 5.8.
The $k$-ary operation $f:B^{k}\to B$ is strictly almost minimal above $\overline{\langle G\rangle}$ if it is almost minimal above $\overline{\langle G\rangle}$ and for every operation $h\in\overline{\langle G\cup\{f\}\rangle}\cap\mathcal{O}^{(k)}$, we have that $h$ generates $f$ together with $G$ in the sense that $f\in\overline{\langle G\cup\{h\}\rangle}\cap\mathcal{O}^{(k)}$.
Strict almost minimality is a strictly weaker condition than minimality since it is possible that an almost minimal $f$ generates a minimal operation of strictly higher arity than itself. Strictness will be obvious when comparing strictly almost minimal $G$-quasi-minorities to minimal $G$-quasi-minorities.
Lemma 5.9.
Let $G\curvearrowright B$ where $G$ is a Boolean group acting freely on $B$. Suppose that $\mathfrak{m}$ is a $G$-quasi-minority strictly almost minimal above $\langle G\rangle$. Fix $a,b\in B$. Then, the function $\mathfrak{m}_{ab}:B\to B$ given by $c\mapsto\mathfrak{m}(a,b,c)$ is injective.
Proof.
From Lemma 5.6, the operation
$\mathfrak{m}(x,y,\mathfrak{m}(x,y,z))$
is essentially unary depending on the variable $z$. Otherwise, it would have to be a genuine orbit/semiprojection. But that would contradict strict almost minimality of $\mathfrak{m}$ since an orbit/semiprojection only generates other orbit/semiprojections by Lemma 2.12. Hence, we have that
$\mathfrak{m}(x,y,\mathfrak{m}(x,y,z))=\mathfrak{m}(x,x,\mathfrak{m}(x,x,z))=\mathfrak{m}(x,x,\mathfrak{m}(z,z,z))=z.$ (12)
In particular,
$\mathfrak{m}(x,y,z)=w\Rightarrow\mathfrak{m}(x,y,w)=z.$ (13)
Suppose that
$\mathfrak{m}_{ab}(c)=\mathfrak{m}_{ab}(c^{\prime})=d.$
From 13, that implies
$\mathfrak{m}(a,b,d)=c\text{ and }\mathfrak{m}(a,b,d)=c^{\prime},$
yielding that $c=c^{\prime}$. Hence, $\mathfrak{m}_{ab}$ is injective. ∎
Lemma 5.10.
Let $G\curvearrowright B$ be a Boolean group acting freely on $B$. Suppose that $\mathfrak{m}$ is a $G$-quasi-minority strictly almost minimal above $\langle G\rangle$. Then, the $4$-ary relation given by
$\mathfrak{m}(x,y,z)=w$
is symmetric.
Proof.
From Lemma 5.6 and the same calculations as equation 12, for any permutation of $\{1,2,3\},\sigma\in S_{3}$,
$\mathfrak{m}(x_{1},x_{2},\mathfrak{m}(x_{\sigma 1},x_{\sigma 2},x_{\sigma 3}))=x_{3}.$
Hence, by injectivity of $\mathfrak{m}_{ab}$, $\mathfrak{m}(x,y,z)$ is a symmetric operation. Finally, from this and 13, we get that the relation $\mathfrak{m}(x,y,z)=w$ is symmetric. ∎
Remark 5.11.
From the above two lemmas one can also deduce that minorities strictly almost minimal above the trivial group are Steiner $3$-quasigroups and the well-known fact [lindner1978steiner] that the latter induce a Steiner quadruple system on the domain.
Lemma 5.12.
Let $G\curvearrowright B$ be a Boolean group acting freely on $B$. Suppose that $\mathfrak{m}$ is a $G$-quasi-minority strictly almost minimal above $\langle G\rangle$. Then for all $\alpha,\beta,\gamma\in G$,
$\mathfrak{m}(\alpha x,\beta y,\gamma z)=\alpha\beta\gamma\mathfrak{m}(x,y,z).$
Proof.
From Lemma 5.6, we know that for $\alpha\in G$,
$\mathfrak{m}(x,y,\mathfrak{m}(\alpha x,\alpha y,z))$
is a projection onto the last coordinate. In particular, since $G$ is Boolean,
$\mathfrak{m}(x,y,\mathfrak{m}(\alpha x,\alpha y,z))\approx z\approx\mathfrak{m}(x,y,\mathfrak{m}(x,y,z)).$
But by Lemma 5.9, this means that
$\mathfrak{m}(\alpha x,\alpha y,z)\approx\mathfrak{m}(x,y,z).$ (14)
More generally, by the same argument, for $\tau_{1},\tau_{2},\tau_{3}\in G$ such that for $i,j,k\in\{1,2,3\}$ distinct $\tau_{i}=\tau_{j}=\alpha$ and $\tau_{k}=1$,
$\mathfrak{m}(\tau_{1}x,\tau_{2}y,\tau_{3}z)\approx\mathfrak{m}(x,y,z).$ (15)
By symmetry (i.e. Lemma 5.10) applied to equation 14,
$\mathfrak{m}(\alpha x,\alpha y,\mathfrak{m}(x,y,z))\approx z.$
However, by equation 15,
$\mathfrak{m}(\alpha x,y,\alpha\mathfrak{m}(x,y,z))\approx z.$
and so, using symmetry again,
$\mathfrak{m}(\alpha x,y,z)\approx\alpha\mathfrak{m}(x,y,z).$
Applying the same reasoning for each variable we obtain the desired equation. ∎
From the preceding Lemmas we obtain:
Corollary 5.13.
Let $G\curvearrowright B$ be a Boolean group acting freely on $B$. Let $\mathfrak{m}$ be a $G$-minority strictly almost minimal above $\langle G\rangle$. Then, $\mathfrak{m}$ is a $G$-invariant Steiner $3$-quasigroup.
Proof.
By assumption, $\mathfrak{m}$ is an idempotent minority. By Lemma 5.10, it is symmetric. By Lemma 5.6, $\mathfrak{m}$ satisfies SQS. By Lemma 5.12 it satisfies the $G$-invariance identity Inv. Hence, it is a $G$-invariant Steiner $3$-quasigroup. ∎
Lemma 5.14.
Let $G\curvearrowright B$ be a Boolean group acting freely on $B$. Let $m$ be a $G$-invariant Steiner $3$-quasigroup. Then, it is strictly almost minimal above $\langle G\rangle$.
Proof.
Firstly, note that a $G$-invariant Steiner $3$-quasigroup is always a $G$-minority. We claim that any operation in $\langle G\cup\{{m}\}\rangle\cap\mathcal{O}^{(3)}$ is either essentially unary or of the form $\eta{m}(x,y,z)$ for $\eta\in G$. We prove this by induction on complexity of terms. The base case is trivial noting that ${m}$ is symmetric. Suppose that $g_{1},g_{2},g_{3}\in\langle G\cup\{{m}\}\rangle\cap\mathcal{O}^{(3)}$ are of complexity $\leq n$ and consider
${m}(g_{1}(x,y,z),g_{2}(x,y,z),g_{3}(x,y,z)).$
If $g_{1}=\eta_{1}{m}$ and $g_{2}=\eta_{2}{m}$,
${m}(\eta_{1}{m}(x,y,z),\eta_{2}{m}(x,y,z),g_{3}(x,y,z))$ $\approx\eta_{1}\eta_{2}{m}(g_{3}(x,y,z),g_{3}(x,y,z),g_{3}(x,y,z))$
$\approx\eta_{1}\eta_{2}\beta g_{3}(x,y,z),$
yielding that ${m}(g_{1}(x,y,z),g_{2}(x,y,z),g_{3}(x,y,z))$ is either essentially unary or again of the form $\eta{m}$. Hence, the only non-trivial case that remains to be considered is that of only one of the $g_{i}$ being of the form $\eta{m}$ whilst the others are essentially unary. For example, say $g_{1}=\alpha x,g_{2}=\beta y$ and $g_{3}=\eta{m}$. Then,
${m}(\alpha x,\beta y,\eta{m}(x,y,z))\approx\alpha\beta\eta{m}(x,y,{m}(x,y,z))\approx\alpha\beta\eta z,$
which is again essentially unary.
This completes the proof that every operation in $\langle G\cup\{{m}\}\rangle\cap\mathcal{O}^{(3)}$ is either essentially unary or of the form $\eta{m}(x,y,z)$. Note that
$\langle G\cup\{{m}\}\rangle\cap\mathcal{O}^{(3)}=\overline{\langle G\cup\{{m}\}\rangle}\cap\mathcal{O}^{(3)},$
since by freeness of $G\curvearrowright B$, any two operations in $\langle G\cup\{{m}\}\rangle\cap\mathcal{O}^{(3)}$ disagree on any set of more than two points. Hence, all operations in $\overline{\langle G\cup\{{m}\}\rangle}\cap\mathcal{O}^{(3)}$ also generate ${m}$, and so ${m}$ is strictly almost minimal as desired. ∎
Remark 5.15.
From the proof of Lemma 5.14, we can see that for $\mathfrak{m}$ a $G$-quasi-minority strictly almost minimal above $\langle G\rangle$,
$\left((\overline{\langle G\cup\{\mathfrak{m}\}\rangle})\setminus\langle G\rangle\right)\cap\mathcal{O}^{(3)}=\{\eta\mathfrak{m}|\eta\in G\}.$
In particular, any non-essentially unary ternary operation generated by $\mathfrak{m}$ together with $G$ is another $G$-quasi-minority. This will be relevant later when considering strictly almost minimal operations above $\langle\mathbb{Z}_{2}\rangle$.
From Corollary 5.13 and Lemma 5.14 we obtain the following:
Theorem 5.16.
Let $G\curvearrowright B$ be a Boolean group acting freely on $B$. Then, $\mathfrak{m}$ is a $G$-minority strictly almost minimal above $\langle G\rangle$ if and only if it is a $G$-invariant Steiner $3$-quasigroup.
Lemma 5.17.
Let $G\curvearrowright B$ be a Boolean group acting freely on $B$. Let $\mathfrak{m}$ be a $G$-minority minimal above $\langle G\rangle$. Then, $\mathfrak{m}$ is a $G$-invariant Boolean Steiner $3$-quasigroup.
Proof.
Since $\mathfrak{m}$ is minimal, and so strictly almost minimal, from Corollary 5.13, it is a $G$-invariant Steiner $3$-quasigroup. Hence, we just need to show that $\mathfrak{m}$ satisfies the Boolean law Bool. To prove this, consider
$C(x,y,z,w):=\mathfrak{m}(x,\mathfrak{m}(x,z,w),\mathfrak{m}(z,y,w)).$
One can compute that $C(x,y,z,w)$ is a $4$-ary orbit/semiprojection such that for all $a,b,c,d\in B$ with at least two of them in the same orbit,
$C(a,b,c,d)=b.$
By minimality of $\mathfrak{m}$, it must be an actual projection. Hence,
$\mathfrak{m}(x,\mathfrak{m}(x,z,w),\mathfrak{m}(z,y,w))\approx y.$ (16)
Since $\mathfrak{m}(x,y,z)=w$ is a symmetric $4$-ary relation (Lemma 5.10), we can switch $y$ and $\mathfrak{m}(x,z,w)$ in the identity 16 to obtain the Boolean law. ∎
Lemma 5.18.
Let $G\curvearrowright B$ be a Boolean group acting freely on $B$. Suppose that $m$ is a $G$-invariant Boolean Steiner $3$-quasigroup. Then,
$\overline{\langle\{m\}\cup G\rangle}=\langle\{m\}\cup G\rangle,$
and every operation $g\in\overline{\langle\{m\}\cup G\rangle}$ depending on all variables is of the form
$g(x_{1},\dots,x_{r})=\alpha\sum_{i\leq r}x_{i},$
where $\alpha\in G$, $r\neq 2$, and $m(x,y,z)=x+y+z$ for a Boolean group $(B,+,e)$
Proof.
Being a Boolean Steiner $3$-quasigroup, $m(x,y,z)=x+y+z$ for a Boolean group $(B,+,e)$. By an induction on complexity, we can see that any $g\in\langle\{m\}\cup G\rangle$ is (up to permuting the $x_{i}$) of the form
$g(x_{1},\dots,x_{k})=\alpha\sum_{i\leq r}x_{i},$
where $\alpha\in G$, and $r\leq k$ with $r\neq 2$ since otherwise, $g$ would be an essential binary operation contradicting almost minimality of $m$ (which we have by Lemma 5.14). This can be seen since the $G$-invariance identity (Inv) allows us to move all automorphisms outside of any term and since $m(x,y,z)$ is of the form $x+y+z$. Note that each such $g$ can be obtained as for $r\geq 3$
$\alpha\sum_{i\leq r}x_{i}=\alpha m(e,x_{1},m(e,x_{2},m(\dots,m(e,x_{r-1},x_{r})\dots))).$
Now, we prove that $\overline{\langle\{m\}\cup G\rangle}=\langle\{m\}\cup G\rangle$. Consider without loss of generality $g_{1},g_{2}\in\langle\{m\}\cup G\rangle$ depending on all variables. We show that they disagree on some finite set. We may assume without loss of generality that $g_{1}$ and $g_{2}$ have the same arity, and so are of the form
$g_{1}(x_{1},\dots,x_{r})=\alpha\sum_{i\leq r}x_{i}\text{ and }g_{2}(x_{1},\dots,x_{r})=\beta\sum_{i\leq r}x_{i},$
for $\alpha\neq\beta\in G$. Then, for $e$ the identity in $(B,+,e)$, since $G\curvearrowright B$ is free we have that
$g_{1}(e,\dots,e)=\alpha e\neq\beta e=g_{2}(e,\dots,e).$
Hence, we can conclude that $\overline{\langle\{m\}\cup G\rangle}=\langle\{m\}\cup G\rangle$, as desired, and the lemma follows. ∎
Lemma 5.19.
Let $G\curvearrowright B$ be a Boolean group acting freely on $B$. Suppose that $m$ is a $G$-invariant Boolean Steiner $3$-quasigroup. Then, $m$ is minimal above $\langle G\rangle$.
Proof.
Consider $g\in\langle\{m\}\cup G\rangle$ of arity $k\geq 4$ depending on all variables. By Lemma 5.18, we know that
$g(x_{1},\dots,x_{k})=\alpha\sum_{i\leq k}x_{i}.$
Then,
$\alpha g(x_{1},x_{2},x_{3},x_{1},\dots,x_{1})\approx m(x_{1},x_{2},x_{3}),$
yielding that $m\in\langle\{g\}\cup G\rangle$. Hence, by Lemma 5.18, strict almost minimality and the fact that any operation $g\in\langle\{m\}\cup G\rangle$ of arity $k\geq 4$ generates $m$ back together with $G$, we can deduce that $m$ is indeed minimal above $\langle G\rangle$. ∎
Remark 5.20.
Note that essentially the same proofs also yields that a Boolean Steiner $3$-quasigroup is minimal above $\langle\mathrm{Id}\rangle$ and that any minority minimal above $\langle\mathrm{Id}\rangle$ is a Boolean Steiner $3$-quasigroup, thus recovering the original description of idempotent minimal minorities of Rosenberg [fivetypes].
Hence, from Lemma 5.17 and Lemma 5.19, we can deduce that minimal $G$-minorities correspond to $G$-invariant Boolean Steiner $3$-quasigroups:
Theorem 5.21.
Let $G\curvearrowright B$ be a Boolean group acting freely on $B$. Then, $\mathfrak{m}$ is a $G$-quasi-minority minimal above $\langle G\rangle$ if and only if it is a $G$-invariant Boolean Steiner $3$-quasigroup.
5.2 Characterising $G$-invariant Steiner $3$-quasigroups
Remark 5.22.
Let $G\curvearrowright B$ be a Boolean group acting freely on $B$. Let $m$ be a $G$-invariant Steiner $3$-quasigroup. Then, $m$ induces a Steiner $3$-quasigroup on the orbits of $G$, $\tilde{m}:\mathrm{Orb}(G)\to\mathrm{Orb}(G)$, where for $\tilde{a},\tilde{b},\tilde{c}\in\mathrm{Orb}(G)$,
$\tilde{m}(\tilde{a},\tilde{b},\tilde{c})=\tilde{d},$
where $\tilde{d}$ is the unique orbit of $G$ such that for some/all $a\in\tilde{a},b\in\tilde{b},c\in\tilde{c}$,
$m(a,b,c)\in\tilde{d}.$
This is a well-defined operation due to the $G$-invariance identity (Inv). It is then easy to show that $m$ is a symmetric idempotent minority satisfying the condition to be a Steiner $3$-quasigroup. Moreover, if $m$ is Boolean, then so is $\tilde{m}$.
Definition 5.23.
Let $G\curvearrowright B$ be a Boolean group acting freely on $B$. Let $X=\{x_{i}|i\in\mathrm{Orb}(G)\}$ be a set of representatives of the orbits of $G$. Let
$E:=\{(\tilde{x},\tilde{y},\tilde{z},\tilde{w})\in\mathrm{Orb}(G)^{4}\ |\ \tilde{m}(\tilde{x},\tilde{y},\tilde{z})=\tilde{w}\}.$
A $\bm{G}$-weighted Steiner $\bm{3}$-quasigroup (with respect to $X$) consists of a triplet $(m,\tilde{m},f)$, where
  • The function $f:E\to G$ is such that
    $f(\tilde{a},\tilde{b},\tilde{c},\tilde{d})=0$ (17)
    whenever |$|\{\tilde{a},\tilde{b},\tilde{c},\tilde{d}\}|<4$;
  • The ternary operation $\tilde{m}:\mathrm{Orb}(G)^{3}\to\mathrm{Orb}(G)$ is a Steiner $3$-quasigroup;
  • The ternary operation $m:B^{3}\to B$ is defined as follows: given $a,b,c\in B$,
    $m(a,b,c)=\alpha\beta\gamma f(\tilde{a},\tilde{b},\tilde{c},\tilde{m}(\tilde{a},\tilde{b},\tilde{c}))x_{\tilde{m}(\tilde{a},\tilde{b},\tilde{c})}\;,$ (18)
    where $a=\alpha x_{\tilde{a}},b=\beta x_{\tilde{b}},c=\gamma x_{\tilde{c}}$.
Note that this is well-defined since the action of $G$ on $B$ is free and so there is a unique $\alpha$ such that $a=\alpha x_{\tilde{a}}$.
Lemma 5.24.
Let $G\curvearrowright B$ be a Boolean group acting freely on $B$. Let $X=\{x_{i}|i\in\mathrm{Orb}(G)\}$ be a set of representatives of the orbits of $G$. Consider a $G$-weighted Steiner $3$-quasigroup $(m,\tilde{m},f)$. Then, $m$ is a $G$-invariant Steiner $3$-quasigroup.
Proof.
From the condition on $f$ given by 17, we can see that $m$ is idempotent. In fact, for $a=\alpha x_{\tilde{a}}$,
$m(a,a,a)=\alpha f(\tilde{a},\tilde{a},\tilde{a},\tilde{m}(\tilde{a},\tilde{a},\tilde{a}))x_{\tilde{a}}=\alpha x_{\tilde{a}}=a,$ (19)
where the first equality is the definition of $m$ in 18 and the second is 17. From the definition 18, we can see that $m$ is a symmetric minority satisfying the $G$-invariance identity (Inv). Hence, we only need to check that $m$ is indeed a Steiner $3$-quasigroup. We verify this below. Let $a=\alpha x_{\tilde{a}},b=\beta x_{\tilde{b}},c=\gamma x_{\tilde{c}}$
$m(a,b,m(a,b,c))$ $=m(a,b,\alpha\beta\gamma f(\tilde{a},\tilde{b},\tilde{c},\tilde{m}(\tilde{a},\tilde{b},\tilde{c}))x_{\tilde{m}(\tilde{a},\tilde{b},\tilde{c})})$ (20)
$=\gamma f(\tilde{a},\tilde{b},\tilde{c},\tilde{m}(\tilde{a},\tilde{b},\tilde{c}))f(\tilde{a},\tilde{b},\tilde{c},\tilde{m}(\tilde{a},\tilde{b},\tilde{c}))x_{\tilde{c}}=c.$ (21)
The first equality is just Definition 18. The second is also definition 18 noting that since $\tilde{m}$ is a Steiner $3$-quasigroup, $\tilde{m}(\tilde{a},\tilde{b}\tilde{m}(\tilde{a},\tilde{b},\tilde{c}))=\tilde{c}$. The final equality follows from $G$ being Boolean and the definition of $c$. ∎
Lemma 5.25.
Let $G\curvearrowright B$ be a Boolean group acting freely on $B$. and let $X=\{x_{i}|i\in\mathrm{Orb}(G)\}$ be a set of representatives of the orbits of $G$. Let $m$ be a $G$-invariant Steiner $3$-quasigroup. Then, $(m,\tilde{m},f)$ is a $G$-weighted Steiner $3$-quasigroup with respect to $X$, where $\tilde{m}$ is the Steiner $3$-quasigroup on the orbits of $G$ induced by $m$ defined in Remark 5.22 and $f$ is the function $f:E\to G$ where, for $m(x_{\tilde{a}},x_{\tilde{b}},x_{\tilde{c}})=\delta x_{\tilde{m}(\tilde{a},\tilde{b},\tilde{c})}$,
$f(\tilde{a},\tilde{b},\tilde{c},\tilde{m}(\tilde{a},\tilde{b},\tilde{c}))=\delta.$
Proof.
We only need to check the identity from 18. For $a=\alpha x_{\tilde{a}},b=\beta x_{\tilde{b}},c=\gamma x_{\tilde{c}}$, and $m(x_{\tilde{a}},x_{\tilde{b}},x_{\tilde{c}})=\delta x_{\tilde{m}(\tilde{a},\tilde{b},\tilde{c})}$,
$m(a,b,c)$ $=\alpha\beta\gamma m(x_{\tilde{a}},x_{\tilde{b}},x_{\tilde{c}})$ (22)
$=\alpha\beta\gamma\delta x_{\tilde{m}(\tilde{a},\tilde{b},\tilde{c})}$ (23)
$=\alpha\beta\gamma f(\tilde{a},\tilde{b},\tilde{c},\tilde{m}(\tilde{a},\tilde{b},\tilde{c}))x_{\tilde{m}(\tilde{a},\tilde{b},\tilde{c})}\;.$ (24)
Here the first equality follows from the $G$-invariance identity (Inv) and the last equality is just from the definition of $f$. ∎
Remark 5.26.
Lemmas 5.24 and 5.25 imply necessary and sufficient conditions for the existence of a $G$-invariant Steiner $3$-quasigroup. In particular, from Hanani [hanani1960quadruple], we know that Steiner quadruple systems on $s$ many vertices exist if and only if $s\equiv 2$ or 4mod $64\ \mathrm{mod}$ 6 (or is infinite). Hence, for $G\curvearrowright B$ a Boolean group acting freely on $B$ with $s$ many orbits, there is a $G$-invariant Steiner $3$-quasigroup if and only if $s\equiv 2$ or 4mod $64\ \mathrm{mod}$ 6 (or is infinite). Indeed, given the lack of restrictions on $f$ in Definition 5.23, we can compute that for $s$-many $G$-orbits the number of $G$-invariant Steiner $3$-quasigroups is
$\mathrm{nSQS}(s)|G|^{\frac{s(s-1)(s-2)}{24}},$
where $\mathrm{nSQS}(s)$ is the number of Steiner quadruple systems on $s$ many vertices and $s(s-1)(s-2)/{24}$ is the number of $4$-hyperedges (a.k.a. blocks) in a Steiner quadruple system of size $s$.
Definition 5.27.
A $G$-weighted Steiner $3$-quasigroup $(m,\tilde{m},f)$ is Boolean if $\tilde{m}$ is Boolean and, for all $\tilde{a},\tilde{b},\tilde{c},\tilde{d},$
$f(\tilde{a},\tilde{b},\tilde{c},\tilde{a}\oplus\tilde{b}\oplus\tilde{c})f(\tilde{a},\tilde{b},\tilde{d},\tilde{a}\oplus\tilde{b}\oplus\tilde{d})=f(\tilde{c},\tilde{d},\tilde{a}\oplus\tilde{b}\oplus\tilde{c},\tilde{a}\oplus\tilde{b}\oplus\tilde{d}),$ (25)
writing $\tilde{m}(x,y,z)=x\oplus y\oplus z$, where $\oplus$ is addition in a Boolean group on $\mathrm{Orb}(G)$.
Lemma 5.28.
Let $G\curvearrowright B$ be a Boolean group acting freely on $B$. Let $X=\{x_{i}|i\in\mathrm{Orb}(G)\}$ be a set of representatives of the orbits of $G$. Let $m$ be a $G$-invariant Boolean Steiner $3$-quasigroup and let $(m,\tilde{m},f)$ be the associated $G$-weighted Steiner $3$-quasigroup from Lemma 5.25. Then, $(m,\tilde{m},f)$ is Boolean.
Remark 5.29.
The condition 25 tells us that in the Steiner quadruple system induced by $\tilde{m}$ on $\mathrm{Orb}(G)$, $f$ assigns $4$-hyperedges elements of $G$ so that whenever two $4$-hyperedges $H_{1}$ and $H_{2}$ intersect in two vertices, the $4$-hyperedge containing the remaining four vertices in $H_{1}\cup H_{2}$ is assigned value $\alpha\beta$ by $f$, were $H_{1}$ and $H_{2}$ were assigned values $\alpha$ and $\beta$ respectively.
Proof of Lemma 5.28.
Consider $x_{\tilde{d}},x_{\tilde{c}},x_{\tilde{a}},x_{\tilde{b}}$. By the Boolean law,
$m(x_{\tilde{d}},x_{\tilde{c}},m(x_{\tilde{a}},x_{\tilde{c}},x_{\tilde{b}}))=m(x_{\tilde{a}},x_{\tilde{b}},x_{\tilde{d}}).$ (26)
But being a $G$-weighted Steiner $3$-quasigroup, we also have the following for the left-hand side of 26:
$m(x_{\tilde{c}},x_{\tilde{d}},m(x_{\tilde{a}},x_{\tilde{b}},x_{\tilde{c}}))$ $=f(\tilde{a},\tilde{b},\tilde{c},\tilde{a}\oplus\tilde{c}\oplus\tilde{b})m(x_{\tilde{d}},x_{\tilde{c}},x_{\tilde{a}\oplus\tilde{c}\oplus\tilde{b}})$ (27)
$=f(\tilde{a},\tilde{b},\tilde{c},\tilde{a}\oplus\tilde{c}\oplus\tilde{b})f(\tilde{c},\tilde{d},\tilde{a}\oplus\tilde{b}\oplus\tilde{d},\tilde{a}\oplus\tilde{b}\oplus\tilde{c})x_{\tilde{a}\oplus\tilde{b}\oplus\tilde{d}}.$ (28)
Meanwhile, for the right hand side of 26, we have
$m(x_{\tilde{a}},x_{\tilde{b}},x_{\tilde{d}})=f({\tilde{a}},{\tilde{b}},{\tilde{d}},\tilde{d}\oplus\tilde{a}\oplus\tilde{b})x_{\tilde{a}\oplus\tilde{b}\oplus\tilde{d}}.$
Hence, comparing the two sides of equation 26 we deduce that
$f(\tilde{a},\tilde{b},\tilde{c},\tilde{a}\oplus\tilde{c}\oplus\tilde{b})f(\tilde{d},\tilde{c},\tilde{d}\oplus\tilde{a}\oplus\tilde{b},\tilde{c}\oplus\tilde{a}\oplus\tilde{b})=f({\tilde{a}},{\tilde{b}},{\tilde{d}},\tilde{d}\oplus\tilde{a}\oplus\tilde{b}),$
and since $G$ is Boolean we also get the desired equation.
Lemma 5.30.
Let $G\curvearrowright B$ be a Boolean group acting freely on $B$. Let $X=\{x_{i}|i\in\mathrm{Orb}(G)\}$ be a set of representatives of the orbits of $G$. Let $(m,\tilde{m},f)$ be a Boolean $G$-weighted Steiner $3$-quasigroup. Then, $m$ is a $G$-invariant Boolean Steiner $3$-quasigroup.
Proof.
We know from Lemma 5.24 that $m$ is a $G$-invariant Steiner $3$-quasigroup, so we only need to prove the Boolean law. We have for $a=\alpha x_{\tilde{a}},b=\beta x_{\tilde{b}},c=\gamma x_{\tilde{c}},d=\delta x_{\tilde{d}}$,
$m(a,b,m(c,b,d))$ $=\gamma\beta\delta f(\tilde{c},\tilde{b},\tilde{d},\tilde{c}\oplus\tilde{b}\oplus\tilde{d})m(a,b,x_{\tilde{c}\oplus\tilde{b}\oplus\tilde{d}})$ (29)
$=\alpha\gamma\delta f(\tilde{c},\tilde{b},\tilde{d},\tilde{c}\oplus\tilde{b}\oplus\tilde{d})m(x_{\tilde{a}},x_{\tilde{b}},x_{\tilde{c}\oplus\tilde{b}\oplus\tilde{d}})$ (30)
$=\alpha\gamma\delta f(\tilde{c},\tilde{b},\tilde{d},\tilde{c}\oplus\tilde{b}\oplus\tilde{d})f(\tilde{a},\tilde{b},\tilde{c}\oplus\tilde{b}\oplus\tilde{d},\tilde{c}\oplus\tilde{a}\oplus\tilde{d})x_{\tilde{c}\oplus\tilde{a}\oplus\tilde{d}}$ (31)
$=\alpha\gamma\delta f(\tilde{c},\tilde{a},\tilde{d},\tilde{c}\oplus\tilde{a}\oplus\tilde{d})x_{\tilde{c}\oplus\tilde{a}\oplus\tilde{d}}$ (32)
$=m(a,c,d).$ (33)
The first three identities and the last one are just unravelling the definition. Equation 25 is used in the fourth equality. ∎
Remark 5.31.
Note that whenever $\mathrm{Orb}(G)$ is the domain of a Boolean group (i.e. $|\mathrm{Orb}(G)|=2^{n}$ or is infinite), we can construct a Boolean $G$-weighted Steiner $3$-quasigroup $(m,\tilde{m},f)$. The construction is as follows: take $\tilde{m}(x,y,z)=x\oplus y\oplus z$, where $\oplus$ is addition in a Boolean group on $\mathrm{Orb}(G)$. Assign to each point of $\tilde{a}\in\mathrm{Orb}(G)$ an element of $G$, $g_{\tilde{a}}$. Thus, for $\tilde{a},\tilde{b},\tilde{c}$ in $\mathrm{Orb}(G)$, let
$f(\tilde{a},\tilde{b},\tilde{c},\tilde{a}\oplus\tilde{b}\oplus\tilde{c})=g_{\tilde{a}}+g_{\tilde{b}}+g_{\tilde{c}}+g_{\tilde{a}\oplus\tilde{b}\oplus\tilde{c}}.$
It is easy to verify that $f$ satisfies the conditions of 25 since $G$ is a Boolean group. For $|\mathrm{Orb}(G)|=2^{n}$ for $n\leq 3$ (with few calculations in the $n=3$ case), all Boolean $G$-weighted Steiner $3$-quasigroups are of this form, though it looks implausible this extends to larger numbers of orbits. We can thus conclude the following:
Corollary 5.32.
Let $G\curvearrowright B$ be a Boolean group acting freely on $B$. There is a one-to-one correspondence between $G$-invariant Boolean Steiner $3$-quasigroups $m$ and Boolean $G$-weighted Steiner $3$-quasigroups $(m,\tilde{m},f)$. In particular, there is a $G$-quasi-minority minimal above $\langle G\rangle$ if and only if $|\mathrm{Orb}(G)|=2^{n}$ for some $n\in\mathbb{N}$ or is infinite.
6 Almost minimal operations above ${\langle\mathbb{Z}_{2}\rangle}$
In this section we classify almost minimal operations in the case of $\mathbb{Z}_{2}$ acting freely on $B$. This result is Theorem 6.7 and two new kinds of almost minimal operations appear: odd majorities (Definition 1.26) and odd Malcev operations (Definition 1.27). In Subsection 6.1, we show these operations cannot be strictly almost minimal (or minimal), thus completing the type-classification part of the proof of Theorem A.
Remark 6.1.
Recall the definition of an odd majority $m$ (Definition 1.26). Note that this makes sense in the context of $\mathbb{Z}_{2}$ acting freely on $B$. Given a triplet from $B$, if three elements are all in the same orbit of $\mathbb{Z}_{2}$, two of them must be equal and so $m$ acts as a majority. However, if a triplet contains two distinct elements in the same orbit and a third element $c$ outside that orbit, $m$ takes the value $m(c,c,c)$ for that triplet.
Lemma 6.2.
Let $G\curvearrowright B$. Let $m$ be a quasi-majority almost minimal above $\overline{\langle G\rangle}$. Then $m$ is an odd majority.
Proof.
By Lemma 2.4 we have that for any $\gamma\in G$, $m(x,\gamma y,z)$ is quasi-Malcev, and in particular
$m(x,\gamma x,y)\approx m(y,y,y)\;.$
By the symmetry of the definition of a quasi-majority the same identity holds for any permutation of the variables on the left side. ∎
Corollary 6.3.
Let $G\curvearrowright B$ be a Boolean group acting freely on $B$. Let $M$ be a quasi-Malcev operation almost minimal above ${\langle G\rangle}$. Suppose that $M$ is not a $G$-quasi-minority. Then, for all $\gamma\in G\setminus\{1\},\ M(x,\gamma y,z)$ is an odd-majority.
Proof.
By the five types theorem 1.21, $M(x,\gamma y,z)$ is either a quasi-semiprojection, or it is quasi-Malcev up to permutation of variables, or it is a quasi-majority. By Lemma 4.3, it is not a quasi-semiprojection. By Lemmas 4.4 and 4.5, if any permutation of the variables of $M(x,\gamma y,z)$ is quasi-Malcev, then $M$ is a $G$-quasi-minority. Since this is not the case by assumption, we have that $M(x,\gamma y,z)$ must be a quasi-majority. By Lemma 6.2, $M(x,\gamma y,z)$ is an odd-majority. ∎
Notation 6.4.
For the rest of this section $\gamma$ denotes the non-identity element in $\mathbb{Z}_{2}$.
Recall Definition 1.27, defining an odd Malcev operation.
Lemma 6.5.
Let $\mathbb{Z}_{2}$ act freely on $B$. Let $\gamma$ denote the non-identity element of $\mathbb{Z}_{2}$. Let $M$ be a quasi-Malcev operation such that $M(x,\gamma y,z)$ is an odd-majority. Then, $M$ is an odd Malcev.
Proof.
By hypothesis, $m(x,y,z)=M(x,\gamma y,z)$ is an odd-majority. Moreover, $m(x,\gamma y,z)=M(x,y,z)$. Firstly, we can obtain identity 3 since
$M(x,y,x)\approx m(x,\gamma y,x)\approx m(x,\gamma x,x)\approx M(x,x,x),$
where the intermediate identity follows from $m$ being an odd-majority, and in particular a quasi-majority. For identity 4, we have
$M(x,\gamma x,y)\approx m(x,x,y)\approx m(x,\gamma x,x)\approx M(x,x,x),$
where we use the fact that $m$ is a majority. Similarly, we can prove that
$M(y,\gamma x,x)\approx M(x,x,x),$
which together with the previous identity yields 4. Finally, since $m(x,\gamma y,z)$ is an odd majority,
$M(x,y,\gamma x)\approx m(x,\gamma y,\gamma x)\approx m(\gamma y,\gamma y,\gamma y)\approx m(\gamma y,y,\gamma y)\approx M(\gamma y,\gamma y,\gamma y).$
The above yields identity 5, concluding the argument. ∎
Lemma 6.6.
Let $\mathbb{Z}_{2}$ act on $B$. Let $f$ be a binary operation almost minimal above ${\langle\mathbb{Z}_{2}\rangle}$. Then, $f$ is a binary orbit/semiprojection.
Proof.
Clearly, $f(x,x)$ is an orbit/semiprojection if and only if $\gamma\circ f$ is; since one of these functions is idempotent, we may assume without loss of generality that $f$ is. Since $f$ is almost minimal, $f(x,\gamma x)=h(x)$ for some $h\in\mathbb{Z}_{2}$. If $h$ is the identity, then whenever $a_{1},a_{2}$ are elements from the same orbit, then $f(a_{1},a_{2})=a_{1}$, proving the lemma. If $h=\gamma$, then in the same situation we always have $f(a_{1},a_{2})=\gamma a_{1}$, again proving the statement. ∎
Theorem 6.7 ($\mathbb{Z}_{2}$ case).
Let $\mathbb{Z}_{2}$ act freely on $B$ with $s$-many orbits (where $s$ is possibly infinite). Let $f$ be an almost minimal operation above ${\langle\mathbb{Z}_{2}\rangle}$. Then, $f$ is of one of the following types:
  1. 1.
    a unary operation;
  2. 2.
    a ternary $G$-quasi-minority;
  3. 3.
    an odd majority;
  4. 4.
    an odd Malcev (up to permutation of variables);
  5. 5.
    a $k$-ary orbit/semiprojection for $2\leq k\leq s$.
Proof.
The theorem follows by the five types theorem 1.21 and the results in this section. If $f$ is a quasi-majority, then it must be an odd majority by Lemma 6.2. Suppose that $f$ is, up to permuting variables, quasi-Malcev. Let $M(x,y,z)$ be the corresponding quasi-Malcev operation. If $M(x,y,z)$ is a $G$-quasi-minority, by symmetry, $f$ is also a $G$-quasi-minority. Otherwise, by Corollary 6.3, $M(x,\gamma y,z)$ is an odd-majority, and so by Lemma 6.5, $M(x,y,z)$ is an odd Malcev. Finally, as usual, if $f$ is a quasi-semiprojection it has to be an orbit/semiprojection by Lemma 2.1. ∎
Remark 6.8.
By Lemma 2.4, Lemma 6.2 and Lemma 6.5, we have the following: if $M(x,y,z)$ is almost minimal above ${\langle\mathbb{Z}_{2}\rangle}$, then $M(x,y,z)$ is an odd Malcev if and only if $M(x,\gamma y,z)$ is an odd majority.
Theorem 6.9.
Let $\mathbb{Z}_{2}$ act freely on $B$. Then, there is an odd majority $m(x,y,z)$ almost minimal above ${\langle\mathbb{Z}_{2}\rangle}$. Moreover, $m(x,\gamma y,z)$ is an odd Malcev almost minimal above ${\langle\mathbb{Z}_{2}\rangle}$.
Proof.
Remark 6.8 already explains that if $m(x,y,z)$ is an odd majority almost minimal above ${\langle\mathbb{Z}_{2}\rangle}$, then, $m(x,\gamma y,z)$ is an odd Malcev almost minimal above ${\langle\mathbb{Z}_{2}\rangle}$. Hence, we just need to construct an odd-majority almost minimal above ${\langle\mathbb{Z}_{2}\rangle}$.
It is easy to check that the defining identities of an odd majority $m$ are consistent when the group is a free action of $\mathbb{Z}_{2}$ (even imposing idempotency on $m$); $m$ can then be extended to triplets from distinct orbits arbitrarily. ∎
6.1 Odd majorities and odd Malcevs cannot be minimal
Lemma 6.10.
Let $\mathbb{Z}_{2}$ act freely on $B$. Suppose that $m$ is an idempotent odd majority almost minimal above $\langle\mathbb{Z}_{2}\rangle$. Then,
$\mathfrak{m}^{\star}(x,y,z):=m(x,\gamma m(x,y,z),m(\gamma x,y,z))$
is a $G$-minority.
Proof.
Firstly, we evaluate $m_{1}(x,y,z):=m(\gamma x,y,z)$ on tuples where two elements are equal:
$m_{1}(x,x,y)$ $:=m(\gamma x,x,y)\approx y,$
$m_{1}(x,y,x)$ $:=m(\gamma x,y,x)\approx y,$
$m_{1}(y,x,x)$ $:=m(\gamma y,x,x)\approx x,$
$m_{1}(x,x,x)$ $:=m(\gamma x,x,x)\approx x.$
Now, we evaluate $\mathfrak{m}^{\star}$ on triplets where two elements are equal
$\mathfrak{m}^{\star}(x,x,y)$ $:=m(x,\gamma x,y)\approx y,$
$\mathfrak{m}^{\star}(x,y,x)$ $:=m(x,\gamma x,y)\approx y,$
$\mathfrak{m}^{\star}(y,x,x)$ $:=m(y,\gamma x,x)\approx y,$
$\mathfrak{m}^{\star}(x,x,x)$ $:=m(x,\gamma x,x)\approx x,$
Since $\mathfrak{m}^{\star}$ is almost minimal (being generated by an almost minimal operation) and it is a minority, it must be a $G$-minority. ∎
Corollary 6.11.
Let $\mathbb{Z}_{2}$ act freely on $B$. There are no strictly almost minimal odd majorities or odd Malcev operations above $\langle\mathbb{Z}_{2}\rangle$.
Proof.
It is sufficient to prove there cannot be any strictly almost minimal idempotent odd majority: any odd Malcev generates an odd majority and any odd majority $m(x,y,z)$ generates an idempotent odd majority since either $m$ or $\gamma m$ is idempotent.
From Lemma 5.14, we know that $m$ generates a $G$-minority $\mathfrak{m}^{\star}$. Since $m$ is strictly almost minimal, $\mathfrak{m}^{\star}$ must also be strictly almost minimal. However, we know from Lemma 5.14 and Remark 5.15 that the only ternary operations in $\langle G\cup\{\mathfrak{m}^{\star}\}\rangle$ not in $\langle G\rangle$ are of the form $\eta\mathfrak{m}^{\star}$ for $\eta\in G$, and so other $G$-quasi-minorities. So
$m\not\in\langle G\cup\{\mathfrak{m}^{\star}\}\rangle,$
implying that $m$ is not strictly almost minimal. ∎
Since we know that odd majorities and odd Malcev operations are never minimal, we can deduce the classification of minimal operations above a Boolean group acting freely on a set:
Theorem 6.12 (minimal operations, Boolean case).
Let $G\curvearrowright B$ be a Boolean group acting freely on $B$ with $s$-many orbits (where $s$ is possibly infinite). Let $f$ be a minimal operation above $\langle G\rangle$. Then, $f$ is of one of the following types:
  1. 1.
    a unary operation;
  2. 2.
    a binary operation;
  3. 3.
    a ternary minority of the form $\alpha\mathfrak{m}$ where $\mathfrak{m}$ is a Boolean $G$-invariant Steiner $3$-quasigroup and $\alpha\in G$;
  4. 4.
    a $k$-ary orbit/semiprojection for $3\leq k\leq s$.
Furthermore, all Boolean $G$-invariant Steiner $3$-quasigroups on $B$ are minimal and they exist if and only if $s=2^{n}$ for some $n\in\mathbb{N}$ or is infinite.
Proof.
Theorems 4.7 and 6.7 classify almost minimal operations above a Boolean group acting freely on $B$. Theorem 5.21 tells us that minimal $G$-minorities have to be Boolean $G$-invariant $3$-quasigroups. All non-idempotent ones can be obtained by pre-composing an idempotent one by some $\alpha\in G$, and Corollary 5.32 gives us necessary and sufficient conditions for the existence of a Boolean $G$-invariant $3$-quasigroup. Theorem 6.11 tells us that there cannot be any strictly almost minimal (and thus no minimal) odd majorities or odd Malcev operations, thus concluding the proof. ∎
7 Pálfy’s Theorem for orbit-semiprojections
In Corollary 5.32, we gave necessary and sufficient conditions for the existence of a minimal $G$-quasi-minority. In this section, we investigate the existence of $k$-ary orbit/semiprojections. For semiprojections, a theorem of Pálfy [palfy1982contributions] yields that there is a $k$-ary semiprojection minimal above the clone of projections for each $2\leq k\leq s$ where $s$ is the size of the domain. Using a similar construction, we prove an analogue for orbit/semiprojections in Theorem 7.10. We mimic the idea of Pálfy of constructing an essential semiprojection as close as possible to a projection in order to have a good understanding of the other operations it generates. With this, we prove that in a finite context orbit/semiprojections exist in each arity $2\leq k\leq s$ where $s$ is the number of $G$-orbits. For $B$ infinite, it sounds plausible that there are group actions $G\curvearrowright B$ with $k$ many orbits and no $k$-ary orbit/semiprojections. Still, with the non-constructive twist of Fact 3.1, our result also holds for $\mathrm{Aut}(B)\curvearrowright B$ being the action of the automorphism group of an $\omega$-categorical structure in a finite relational language.
Notation 7.1.
For $a,b\in B$, we write $a\sim b$ to say that $a$ and $b$ are in the same orbit.
Definition 7.2.
Let $G\curvearrowright B$ with $s$ many orbits. Let $b\in B$ be in a $G$-orbit consisting of more than one element. Let $\alpha,\beta\in G$. For finite $k\leq s$, we define the $k$-ary orbit/semiprojection $f^{(\alpha,\beta)}$ as follows: for $(a_{1},\dots,a_{k})\in B^{k}$,
$f^{(\alpha,\beta)}(a_{1},\dots,a_{k}):=\begin{cases}\beta b\text{ if }a_{1}\sim b\text{ and }a_{1},\dots,a_{k}\text{ are all in distinct orbits};\\ \alpha a_{1}\text{ otherwise.}\end{cases}$ (34)
We write $f$ for $f^{(e,e)}$, where $e$ is the identity in $G$.
Remark 7.3.
Note that any $f^{(\alpha,\beta)}$ is in $\langle G\cup\{f\}\rangle$, being obtained by
$\beta f(\beta^{-1}\alpha x_{1},\dots,x_{k}).$
Conversely, $f\in\langle G\cup\{f^{(\alpha,\beta)}\}\rangle$. Also note that for $\gamma\in G$,
$\gamma f^{(\alpha,\beta)}=f^{(\gamma\alpha,\gamma\beta)}.$
Notation 7.4.
For $t$ a term in $\langle G\cup\{f\}\rangle$, we denote by $l(t)$ the leftmost $i$ in some $x_{i}$ appearing in $t$. So, for example
$l(\alpha f(f(\beta x_{2},x_{1},f(x_{3},\gamma x_{1},x_{1})),f(x_{1},x_{2},x_{1}),x_{1}))=2.$
Remark 7.5.
Note that by construction of $f$, for any $t\in\langle G\cup\{f\}\rangle$, we have that for all $(a_{1},\dots,a_{r})\in B^{r}$, $t(a_{1},\dots,a_{r})\sim a_{l(t)}$. This can be easily seen by induction on complexity.
Lemma 7.6.
Let $G\curvearrowright B$ and $f^{(\alpha,\beta)}$ be as in Definition 7.2. Then, for any operations $t_{2},\dots,t_{k}\in\langle G\cup\{f\}\rangle$ and $\alpha,\beta,\gamma,\delta\in G$, we have the following identities
  • If $l(t_{i})=i$ for all $2\leq i\leq k$,
    $f^{(\alpha,\beta)}(\gamma x_{1},t_{2},\dots,t_{k})\approx f^{(\alpha\gamma,\beta)};$ (35)
  • If $l(t_{i})=1$ or $l(t_{i})=l(t_{j})$ for some $2\leq i<j\leq k$,
    $f^{(\alpha,\beta)}(\gamma x_{1},t_{2},\dots,t_{k})\approx\alpha\gamma x_{1};$ (36)
  • If $l(t_{i})=i$ for all $2\leq i\leq k$,
    $f^{(\alpha,\beta)}(f^{(\gamma,\delta)},t_{2},\dots,t_{k})\approx f^{(\alpha\gamma,\beta)};$ (37)
  • If $l(t_{i})=1$ or $l(t_{i})=l(t_{j})$ for some $2\leq i<j\leq k$,
    $f^{(\alpha,\beta)}(f^{(\gamma,\delta)},t_{2},\dots,t_{k})\approx\alpha f^{(\gamma,\delta)}.$ (38)
Proof.
The identities follow directly from Definition 7.2 and Remark 7.5. For example, for 35, given $(a_{1},\dots,a_{k})$, by Remark 7.5, we have that
$f^{(\alpha,\beta)}(\gamma x_{1},t_{2},\dots,t_{k})(a_{1},\dots,a_{k})=f^{(\alpha,\beta)}(\gamma a_{1},a_{2}^{\prime},\dots,a_{k}^{\prime}),$
where $a_{i}^{\prime}\sim a_{i}$. Hence, if $(a_{1},\dots,a_{k})$ were all in distinct orbits and $a_{1}\sim b$, we get that
$f^{(\alpha,\beta)}(\gamma a_{1},a_{2}^{\prime},\dots,a_{k}^{\prime})=\beta b.$
Otherwise, we have
$f^{(\alpha,\beta)}(\gamma a_{1},a_{2}^{\prime},\dots,a_{k}^{\prime})=\alpha\gamma a_{1}.$
This yields the desired identity in 35. For identity 36, by Remark 7.5, we get that
$f^{(\alpha,\beta)}(\gamma x_{1},t_{2},\dots,t_{k})(a_{1},\dots,a_{k})=f^{(\alpha,\beta)}(\gamma a_{1},a_{l(t_{2})}^{\prime},\dots,a_{l(t_{k})}^{\prime}),$
where $a_{l(t_{k})}^{\prime}\sim a_{l(t_{k})}$. In particular, since either $l(t_{i})=1$ or $l(t_{i})=l(t_{j})$ for $2\leq i<j\leq k$, we are never in the first condition of Definition 7.2, and so the identity 36 holds. The final two identities 37 and 38 follow by the same reasoning as the previous two. ∎
Lemma 7.7.
Let $t\in\langle G\cup\{f\}\rangle\setminus\langle G\rangle$. Then, $t$ is of the form $f^{(\alpha,\beta)}$ for some $\alpha,\beta\in G$ up to permuting variables and adding dummy variables.
Proof.
We prove this by induction on complexity of $t$, where the base case is trivial. Now, consider $t\in\langle G\cup\{f\}\rangle\setminus\langle G\rangle$. We may assume without loss of generality that $t$ is of the form $f^{(\alpha,\beta)}(t_{1},\dots,t_{k})$. By inductive hypothesis, we may suitably permute variables to obtain $t^{\prime}$ of the form $f^{(\alpha,\beta)}(t^{\prime}_{1},\dots,t^{\prime}_{k})$, where $t^{\prime}_{1}$ is either of the form $\gamma x_{1}$ or of the form $f^{(\gamma,\delta)}(x_{1},\dots,x_{k})$ for some $\gamma,\delta\in G$. The identities in Lemma 7.6 then tell us that $t^{\prime}$ is either in $\langle G\rangle$ or again of the form $f^{(\alpha^{\prime},\beta^{\prime})}$ for $\alpha^{\prime},\beta^{\prime}\in G$. ∎
Corollary 7.8.
Let $G\curvearrowright B$ with $s$-many orbits. Then, for each finite $1<k\leq s$ there is a $k$-ary orbit/semiprojection $f$ depending on all variables such that for every $h\in\langle G\cup\{f\}\rangle\setminus\langle G\rangle$, $f\in\langle G\cup\{h\}\rangle\setminus\langle G\rangle$ and every $h\in\langle G\cup\{f\}\rangle\setminus\langle G\rangle$ has arity $\geq k$.
Proof.
Use $f$ from Definition 7.2. The statement holds by Lemma 7.7 and Remark 7.3. ∎
Remark 7.9.
Corollary 7.8 yields that for every $G\curvearrowright B$ with $s$-many orbits and $1<k\leq s$, there is a $k$-ary orbit/semiprojection $f$ such that $\langle G\cup\{f\}\rangle$ is minimal in the monoidal interval of all (i.e. not just the closed ones) clones above $\langle G\rangle$. That is, $f$ is minimal above $\langle G\rangle$ in the non-topological sense of minimality. Thus, for minimal operations above permutation groups in the lattice of all clones, we get a version of Theorem A where $k$-ary orbit/semiprojections always exist as minimal above any $G\curvearrowright B$ for all $1<k\leq s$ (cf. Remark 1.29).
Theorem 7.10.
Let $G\curvearrowright B$ with $s$-many orbits with $B$ finite or of the form $\mathrm{Aut}(B)\curvearrowright B$ for $B$ $\omega$-categorical in a finite language. Then, for all finite $1<k\leq s$, there is a $k$-ary orbit/semiprojection minimal above $\overline{\langle G\rangle}$.
Proof.
Use $f$ from Definition 7.2. If $B$ is finite, Corollary 7.8 yields that $f$ is minimal. In the oligomorphic case, by Fact 3.1, $\overline{\langle G\cup\{f\}\rangle}$ contains a minimal operation $t$ above $\overline{\langle G\rangle}$. By Lemma 7.7, $t$ is essentially $k$-ary, and by Lemma 2.12 another orbit/semiprojection, yielding the desired result. ∎
Problem 7.11.
Give an example of $G\curvearrowright B$ with three orbits such that there is no ternary orbit/semiprojection minimal above $\overline{\langle G\rangle}$.
It sounds plausible such a $G$ should exist. However, this does not sound easy to prove: for specific choices of $f$, one can show that there is no orbit/semiprojection in $\overline{\langle G\cup\{f\}\rangle}$ minimal above $\overline{\langle G\rangle}$. However, it is unclear how on an infinite domain one may run an argument that any possible orbit/semiprojection will not be minimal.
8 Solvability in Datalog without binary injections
From Theorem 3.14, we know that whenever $B$ is an $\omega$-categorical model-complete core whose $\mathrm{CSP}$ is not $\mathrm{NP}$-hard, $\mathrm{Pol}(B)$ has a binary essential polymorphism. Frequently in the literature on infinite-domain $\mathrm{CSP}$s, the existence of a binary essential polymorphism is used to deduce the existence of a binary injective polymorphism under some additional tame assumptions [BodKara, bodirsky2010complexity, bodirsky2014minimal, Smoothapp, bodirsky2020complexity]. Binary injective polymorphisms are very helpful to reduce $\mathrm{CSP}(B)$ to its injective variant [Smoothapp], and more generally in most complexity classifications on an infinite domain [bodirsky2010complexity, bodirsky2014minimal, Smoothapp]. Moreover, they naturally appear as one of the Nelson-Opper criteria [nelson1979simplification] for deducing that the generic combination of two $\omega$-structures with no algebraicity and trivial $\mathrm{CSP}$ also has trivial $\mathrm{CSP}$ [bodirsky2020complexity, bodirsky2022tractable].
We conclude by giving a counterexample to the following question of Bodirsky on binary injective polymorphism:
Question 3.
(Question 14.2.6 (27) in [BodCSP]) Does every $\omega$-categorical structure without algebraicity that can be solved by Datalog also have a binary injective polymorphism?
We give a counterexample to this question (Definition 8.1) which is a reduct of two infinite disjoint unary predicates partitioning a countable domain. Datalog programs capture many standard algorithms used in the context of infinite-domain $\mathrm{CSP}$s, such as path consistency (and more generally, $\mathrm{CSP}$s solved by a Datalog program are solvable in polynomial time). We refer the reader to [bodirsky2013datalog] for a discussion of Datalog in an $\omega$-categorical setting and for the definition of a Datalog program. Its significance for Question 3 is the idea that binary injective polymorphisms would always appear for $\omega$-categorical structures with easy enough $\mathrm{CSP}$s. Without the restriction to Datalog, some first-order reducts of $(\mathbb{Q},<)$ have $\mathrm{CSP}$ solvable in polynomial time (though not in Datalog) but have no binary injective polymorphisms (cf. Theorem 1.4 in [bodirsky2022tractable] and Claim 4.8 in [bodirsky2022descriptive]). Interestingly, our counterexample is a reduct of a unary structure and $\mathrm{CSP}$s for such structures have been studied in detail with a complexity dichotomy proved in [bodirsky2018dichotomy, Theorems 1.1 & 1.2] and a descriptive complexity analysis given in [mottet2021symmetries, Theorem 5.1]. We provide a direct proof of solvability in Datalog since the algorithm is particularly simple.
Definition 8.1.
Let $D$ be the first-order structure with domain the disjoint union of two copies of $\mathbb{N}$ in the language with two unary predicates $\{P_{0},P_{1}\}$, where each predicates names one of the two disjoint copies (which we also call $P_{0}$ and $P_{1}$). We define the following relations on $D$:
$r(x,y,z)$ $:=(P_{1}(x)\wedge P_{1}(y))\rightarrow P_{1}(z);$
$p(x,y)$ $:=(x=y\wedge P_{0}(x))\vee(P_{1}(x)\wedge P_{1}(y));$
$q(x,y)$ $:=(x=y\wedge P_{1}(x))\vee(P_{0}(x)\wedge P_{0}(y)).$
Consider the first-order structure
$D^{*}:=(D,P_{0},P_{1},r,p,q).$
Lemma 8.2.
Let $g$ be a binary polymorphism of $D^{*}$. Then, whether $g(a,b)$ is in $P_{0}$ or $P_{1}$ is entirely determined by which $P_{i}$ hold of $a$ and $b$.
Remark 8.3.
This is really saying that $g$ is $\mathrm{Aut}(D)$-canonical.
Proof.
We use subscripts 00 and $1$ to keep track of which vertices are in $P_{0}$ and which are in $P_{1}$. Suppose that for $a_{0},e_{0}\in P_{0}$ and $b_{1},f_{1}\in P_{1}$, we have that $g(a_{0},b_{1})\in P_{1}$ and $g(e_{0},f_{1})\in P_{0}$. Then, take $c_{1},d_{1}\in P_{1}$. Now, consider
$g\begin{pmatrix}a_{0},b_{1}\\ c_{1},d_{1}\\ e_{0},f_{1}\end{pmatrix}=\begin{pmatrix}g(a_{0},b_{1})\\ g(c_{1},d_{1})\\ g(e_{0},f_{1})\end{pmatrix}.$
Since $g$ preserves $P_{1}$, $g(c_{1},d_{1})\in P_{1}$. Note that $r(a_{0},c_{1},e_{0})$ and $r(b_{1},d_{1},f_{1})$. But
$\neg r((g(a_{0},b_{1}),g(c_{1},d_{1}),g(e_{0},f_{1})).$
Since $g$ is a polymorphism of $D^{*}$, this contradicts the initial assumption. We can run the same argument to show that we also can never have $g(b_{1},a_{0})\in P_{1},g(f_{1},e_{0})\in P_{0}$. ∎
Lemma 8.4.
Let $g$ be a binary polymorphism of $D^{*}$. Then, we cannot have, for $a_{0},d_{0}\in P_{0}$ and $b_{1},c_{1}\in P_{1}$,
$g(a_{0},b_{1})\in P_{1}\text{ and }g(c_{1},d_{0})\in P_{1}.$
Proof.
Take $e_{0},f_{0}\in P_{0}$. Then,
$g\begin{pmatrix}a_{0},b_{1}\\ c_{1},d_{0}\\ e_{0},f_{0}\end{pmatrix}=\begin{pmatrix}g(a_{0},b_{1})\\ g(c_{1},d_{0})\\ g(e_{0},f_{0})\end{pmatrix}.$
Since $g$ preserves $P_{0}$, $g(e_{0},f_{0})\in P_{0}$. But then, $g$ does not preserve $r$. ∎
Lemma 8.5.
Suppose that $g$ is a binary polymorphism of $D^{*}$. Then, $g$ is not injective.
Proof.
We need to consider two possibilities. First, suppose that for $a_{0},d_{0}\in P_{0}$ and $b_{1},c_{1}\in P_{1}$,
$g(a_{0},b_{1})\in P_{0}\text{ and }g(c_{1},d_{0})\in P_{1}.$
Now,
$g\begin{pmatrix}a_{0},b_{1}\\ d_{0},b_{1}\\ \end{pmatrix}=\begin{pmatrix}g(a_{0},b_{1})\\ g(d_{0},b_{1})\ \end{pmatrix}.$
Note that $q(a_{0},d_{0})$ and $q(b_{1},b_{1})$. Now, by Lemma 8.2, $g(a_{0},b_{1})\in P_{0}$ and $g(d_{0},b_{1})\in P_{0}$. But since $g$ preserves, $q$, we must have $g(a_{0},b_{1})=g(d_{0},b_{1})$, which breaks injectivity. By Lemma 8.4 and symmetry, the only other case we need to consider is
$g(a_{0},b_{1})\in P_{0}\text{ and }g(c_{1},d_{0})\in P_{0}.$
In this case,
$g\begin{pmatrix}a_{0},b_{1}\\ a_{0},c_{1}\\ \end{pmatrix}=\begin{pmatrix}g(a_{0},b_{1})\\ g(a_{0},c_{1})\end{pmatrix}.$
Now, $p(a_{0},a_{0})$ and $p(b_{1},c_{1})$ hold. However, since $g(a_{0},b_{1}),g(a_{0},c_{1})\in P_{0}$, we must have that $g(a_{0},b_{1})=g(a_{0},c_{1})$, again, breaking injectivity. ∎
Lemma 8.6.
$\mathrm{CSP}(D^{*})$ is solvable in Datalog.
Proof.
Note that the model-complete core of $D$ is a structure $B^{*}$ on the Boolean domain $\{0,1\}$. The relations $p$ and $q$ collapse to the same relation $\{(0,0),(1,1)\}$. Note that $B^{*}$ is preserved by $\mathrm{min}$ and so is solvable in Datalog. ∎
Definition 8.7.
Let $D^{\prime}$ be the structure given by $(D^{*},\neq)$.
Lemma 8.8.
The structure $D^{\prime}$ is a model-complete core.
Proof.
We show that $\overline{\mathrm{Aut}(D^{\prime})}=\mathrm{End}(D^{\prime})$, which is given by
$\mathcal{G}:=\{g_{0}\cup g_{1}|g_{0}\in\mathrm{Inj}(P_{0}),g_{1}\in\mathrm{Inj}(P_{1})\}.$
Let $g\in\mathrm{End}(D)$. We know that $g$ preserves $P_{0}$ and $P_{1}$ and that on each set it must act as an injection since it preserves $\neq$. In particular, $g\in\mathcal{G}$. Conversely, any function in $\mathcal{G}$ preserves all relations of $D^{\prime}$. Finally, note that $\mathrm{Aut}(D^{\prime})=\mathrm{Aut}(D)$ and that the latter is
$\{\sigma_{0}\cup\sigma_{1}|\sigma_{0}\in\mathrm{Bij}(P_{0}),\sigma_{1}\in\mathrm{Bij}(P_{1})\},$
where $\mathrm{Bij}(P_{i})$ is the set of bijections of the countably infinite set denoted by $P_{i}$. Hence, $\overline{\mathrm{Aut}(D^{\prime})}=\mathrm{End}(D^{\prime})$. ∎
Theorem 8.9.
The problem $\mathrm{CSP}(D^{\prime})$ is solvable by a Datalog program.
Head Body
$\mathrm{tc}_{0}(x)$ $P_{0}(x)$
$\mathrm{tc}_{0}(x)$ $\mathrm{tc}_{0}(y),p(x,y)$
$\mathrm{tc}_{0}(x)$ $\mathrm{tc}_{0}(y),p(y,x)$
$\mathrm{tc}_{1}(x)$ $P_{1}(x)$
$\mathrm{tc}_{1}(x)$ $\mathrm{tc}_{1}(y),q(x,y)$
$\mathrm{tc}_{1}(x)$ $\mathrm{tc}_{1}(y),q(y,x)$
$\mathrm{tc}_{1}(x)$ $\mathrm{tc}_{1}(y),\mathrm{tc}_{1}(z),r(y,z,x)$
$\mathrm{tc}_{1}(x)$ $\mathrm{tc}_{1}(y),\mathrm{tc}_{1}(z),r(z,y,x)$
$E(x,y)$ $\mathrm{tc}_{0}(y),p(x,y)$
$E(x,y)$ $\mathrm{tc}_{0}(y),p(y,x)$
$E(x,y)$ $\mathrm{tc}_{1}(y),q(x,y)$
$E(x,y)$ $\mathrm{tc}_{1}(y),q(y,x)$
$E(x,y)$ $E(y,x)$
$E(x,y)$ $E(x,u),E(u,y)$
false $E(x,y),x\neq y$
false $\mathrm{tc}_{0}(x),\mathrm{tc}_{1}(x)$
Table 2: We give a semantic description of the Datalog program. Given an instance $B$ of $\mathrm{CSP}(D^{\prime})$, we add two predicates $\mathrm{tc}_{0}(x),\mathrm{tc}_{1}(x)$ to compute the transitive closure of the predicates $P_{0}$ and $P_{1}$ respectively (that is, we label by $\mathrm{tc}_{i}(x)$ the vertices which are forced to be in $P_{i}$ given the rules for $p,q,$ and $r$). Then, we add a binary relation $E$ to denote the pairs which must be identified in any potential homomorphism $B\to D^{\prime}$ given the definitions of $p$ and $q$ in $D^{\prime}$. Finally, the program returns false if either of two things happen: (1) two vertices which are supposed to remain distinct in the map are actually identified by $E$, (2) a vertex is forced to be both in $P_{0}$ and in $P_{1}$.
Proof.
We give an explicit Datalog program in Table 2 with a semantic description of how the program behaves in the description of the Table. For an instance $B$ of $\mathrm{CSP}(D^{\prime})$, it is clear that if the given Datalog program returns ’false’, $B$ cannot map homomorphically to $D^{\prime}$. So, suppose that our Datalog program returns ’true’. Then, we can label the vertices of $B$ by $\mathrm{tc}_{0},\mathrm{tc}_{1}$, and pairs from $B$ by $\neq$ and by $E$ so that no vertex is labelled by both $\mathrm{tc}_{0}$ and $\mathrm{tc}_{1}$ and so that no pair is labelled by both $E$ and $\neq$. Note that if a vertex in a connected component of $E$ is labelled with $\mathrm{tc}_{i}$, so are all vertices in that component and that the $E$-connected components are complete $E$-graphs. We build a homomorphism $B\to D^{\prime}$ as follows: every vertex in $\mathrm{tc}_{i}$ is sent to a vertex in $P_{i}$. We identify each $E$-connected component with a single vertex and if the vertices in that component are labelled by $\mathrm{tc}_{i}$, we send them to a vertex in $P_{i}$. We don’t make unnecessary identifications (i.e. we do not identify different connected components with the same vertex). Vertices which are not labelled by either $\mathrm{tc}_{i}$ are sent to $P_{1}$. The homomorphism we built preserves $\neq,P_{0}$, and $P_{1}$. It also respects $p,q,$ and $r$ given the rules used by the Datalog program to label vertices with $\mathrm{tc}_{0},\mathrm{tc}_{1}$, and $E$. ∎
Corollary 8.10.
There is a homogeneous finitely bounded model-complete core without algebraicity whose $\mathrm{CSP}$ can be solved by a Datalog program but which has no binary injective polymorphism.
Proof.
The structure $D$ is clearly homogeneous and finitely bounded without algebraicity and so is $D^{\prime}$. We know $D^{\prime}$ is a model-complete core due to Lemma 8.8. We know that $D^{\prime}$ is solvable by a Datalog program by Lemma 2. Finally, we know that $D^{\prime}$ has no binary injections from Lemma 8.5. ∎
[…]